### v1

# Imaginarity of quantum channels: Refinement and Alternative

Abstract
At present, quantum channels have been widely concerned, and many ways to quantify quantum channels have been proposed, which has led to the generation of many resource theories for quantum channels. We add strong monotonicity and convexity to the requirement of imaginarity measure of quantum channels to make the measure proper. We also introduce an alternative framework to simplify the process of verifying whether a quantifier is a proper measure. We present three imaginarity measures of quantum channels via on the robustness, the trace norm and entropy, respectively. Some properties are also given.
Keywords: quantum resource theories, quantum channels, imaginarity measures, alternative framework
1 Introduction
Imaginary numbers provide a broad perspective and additional methods for modern scientific research. Quantum resource theory can establish new connections among various research directions. With the development of quantum resource theory, the imaginarity resource theory [1] proposed in recent years has attracted increasing attention. Some examples of applications, the imaginarity plays a crucial role in state discrimination [2], the imaginarity operation is more advantageous to distinguish the quantum channel without the help of the auxiliary system [3]. Many imaginarity measures have been found for quantum states, the fidelity [2], the convex roof [4], the geometric [3] and some imaginarity measures induced by entropies [5, 6], some of which can establish relations due to their own mathematical properties [6, 7].
Just like quantum states, quantum channels also deserve attention. Some studies begin to investigate the properties of quantum channels [8, 9, 10, 11, 12]. Resource theories of quantum channels was established in [13], and many resource theories of quantum channels have been proposed, such as coherence of quantum channels [14], entanglement of quantum channels [15] and so on [16, 17]. Based on this, more measures for quantum channels have been found [18, 19, 20, 21].
In order to simplify the process of verifying whether a quantifier can be a coherence measure, alternative framework for quantifying coherence have been proposed [22] in the coherence resource theory. Similar alternative frameworks can be established in many other resource theories, such as imaginarity resource theory [7], coherence of quantum channels [23].
The concept of the imaginarity of quantum channels was first introduced in [24], and the definitions of free channels and free superchannels have been visited again recently [25]. It can be seen that the concept is worthing considing. Similar to the coherence of quantum channels, in this paper, we refine the imaginarity of quantum channels and introduce alternative framework to facilitate the verification of whether a quantifier can be a suitable measure. We also show three imaginarity measures of quantum channels. Some properties of the measure are discussed.
2 Refinement for the imaginarity of quantum channels
Let $H_{A}$ and $H_{B}$ be two Hilbert space with dimensions $|A|$ and $|B|$, and orthonormal basis $\left\{|j\rangle\right\}_{j=0}^{|A|-1}=\left\{|k\rangle\right\}_{k=0}^{|A|-1}$ and $\left\{|\alpha\rangle\right\}_{\alpha=0}^{|B|-1}=\left\{|\beta\rangle\right\}_{\beta=0}^{|B|-1}$. We assume $|a\rangle\langle a|=E_{a,a}$ without loss of generality, that is, it happens to be the a-th diagonal. Denote $\mathcal{D}_{A}$ and $\mathcal{D}_{B}$ be the set of all density operators on $H_{A}$ and $H_{B}$, and $\mathcal{C}_{AB}$ be the set of all channels from $\mathcal{D}_{A}$ to $\mathcal{D}_{B}$. A quantum channel $\phi\in\mathcal{C}_{AB}$ can be represented by Choi matrix,
$J_{\phi}=\sum_{j,k}|j\rangle\langle k|\otimes\phi(|j\rangle\langle k|)=\sum_{j,k,\alpha,\beta}\phi_{j,k,\alpha,\beta}|j\rangle\langle k|\otimes|\alpha\rangle\langle\beta|$
with $\phi_{j,k,\alpha,\beta}=\langle\alpha|\phi(|j\rangle\langle k|)|\beta\rangle$.
The Choi state of $\phi$ is $\frac{J_{\phi}}{|A|}\triangleq\chi_{\phi}$. We call channel $\phi$ a real channel if $\chi_{\phi}$ is an real matrix, and denote $\mathcal{RC}_{AB}$ the set of all the real channels. The free channel is defined as the real channel. This is exactlt the free operation in imaginarity resource theory, thus it is appropriate to define the free channel by Choi state.
Let $\mathcal{SC}_{ABA^{\prime}B^{\prime}}$ be the sets of all superchannels from $\mathcal{C}_{AB}$ to $\mathcal{C}_{A^{\prime}B^{\prime}}$. It can be known that [14, 26] superchannel $\Theta$ also has the expression of Kraus operators $\left\{M_{m}\right\}_{m}$ with $\chi_{\Theta(\phi)}=\sum_{m}M_{m}\chi_{\phi}M_{m}^{\dagger}\triangleq\tilde{\Theta}(\chi_{\phi})$, where $\tilde{\Theta}$ is the channel corresponding to the superchannel. We will not distinguish the signs, $\Theta$ or $\tilde{\Theta}$, since it can be judged by the context.
We call a superchannel $\Theta:\mathcal{C}_{AB}\longrightarrow\mathcal{C}_{A^{\prime}B^{\prime}}$ a real superchannels if $\Theta(\phi)\in\mathcal{RC}_{A^{\prime}B^{\prime}}$ for any channels $\phi\in\mathcal{RC}_{AB}$. We denote the set of all the real superchannels by $\mathcal{RSC}_{ABA^{\prime}B^{\prime}}$. The free superchannel is defined as the real superchannel.
In [24], the author presented a framework for quantifying the imaginarity of quantum channels, they introduced the imaginarity measure $C(\phi)\triangleq C(\chi_{\phi})$ of quantum channels should satisfy the following conditions:
$\bm{(C_{1})}$. Faithfulness: $C(\phi)\geqslant 0$, and $C(\phi)=0$ if and only if $\phi\in\mathcal{RC}_{AB}$;
$\bm{(C_{2})}$. Monotonicity: $C(\phi)\geqslant C(\Theta(\phi))$ for any real superchannel $\Theta\in\mathcal{RSC}_{ABA^{\prime}B^{\prime}}$.
Strong monotonicity and convexity often have a great use and a proper quantum resource theory should include them. So we add these to the conditions above, and then the imaginarity measure $C$ of quantum channels becomes the more familiar and complete.
$\bm{(C_{3})}$. Strong monotonicity: Let $\Theta\in\mathcal{RSC}_{ABA^{\prime}B^{\prime}}$ with $\Theta(\cdot)=\sum_{m}M_{m}\cdot M_{m}^{\dagger}$, then $C(\phi)\geqslant\sum_{m}p_{m}C(\phi_{m})$ where $p_{m}=\text{Tr}(M_{m}\chi_{\phi}M_{m}^{\dagger})$ and $\chi_{\phi_{m}}=M_{m}\chi_{\phi}M_{m}^{\dagger}/p_{m}$
$\bm{(C_{4})}$. Convexity: For any $\left\{\phi_{m}\right\}\subset\mathcal{C}_{AB}$ and any probability distribution $\left\{p_{m}\right\}$, $C(\sum_{m}p_{m}\phi_{m})\leqslant\sum_{m}p_{m}C(\phi_{m})$.
We consider that a proper imaginarity measure of quantum channels should satisfy the condition $\bm{(C_{1})}$$\bm{(C_{4})}$. In this way, we can obtain some properties of the imaginarity measure $C$.
Property 1.
Monotonicity under composition with any real channels, that is,
$(a)$. Left composition: for any $\psi\in\mathcal{RC}_{BD}$ and any $\phi\in\mathcal{C}_{AB}$, we have: $C(\psi\circ\phi)\leqslant C(\phi)$.
$(b)$. Right composition: for any $\psi\in\mathcal{RC}_{AB}$ and any $\phi\in\mathcal{C}_{BD}$, we have: $C(\phi\circ\psi)\leqslant C(\phi)$.
Proof.
We only prove $(a)$, because $(b)$ is analogous.
For $\psi\in\mathcal{RC}_{AB}$ and $\phi\in\mathcal{C}_{BD}$, we consider a superchannel $\Theta_{\psi}$ with $\Theta_{\psi}=\psi\circ\phi$. Due to the condition $\bm{(C_{2})}$, we try to explain $\Theta\in\mathcal{RSC}_{ABAD}$. In fact, when superchannel $\Theta_{\psi}$ acts on any real channels, it is impossible to generate imaginarity resource since the channel $\psi$ is real. We can know that the superchannel $\Theta_{\psi}$ is a real superchannel from the definition. ∎
We can also see that if $\psi$ is invertible, $C(\psi\circ\phi)=C(\phi)$ holds.
Property 2.
The combination of imaginarity measures is still an imaginarity measure, that is, for any $\left\{C_{j}\right\}$ and any $\left\{q_{j}\right\}$ with $q_{j}\geqslant 0$, $C\triangleq\sum_{j}q_{j}C_{j}$ is an imaginarity measure .
Proof.
It is obvious that $C(\phi)\geqslant 0$ and due to $q_{j}\geqslant 0$, $C(\phi)=0$ is equivalent to every $C_{j}(\phi)=0$, certainly it indicate $\phi\in\mathcal{RC}$. For any real superchannel $\Theta$, we have $C(\phi)=\sum_{j}q_{j}C_{j}(\phi)\geqslant\sum_{j}q_{j}C_{j}(\Theta(\phi))=C(\Theta(\phi))$.
For $\Theta\in\mathcal{RSC}$ with $\Theta(\cdot)=\sum_{m}M_{m}\cdot M_{m}^{\dagger}$, the strong monotonicity holds:
$C(\phi)=\sum_{j}q_{j}C_{j}(\phi)\geqslant\sum_{j}q_{j}\sum_{m}p_{m}C_{j}(\phi_{m})=\sum_{m}p_{m}\sum_{j}q_{j}C_{j}(\phi_{m})=\sum_{m}p_{m}C(\phi_{m}).$
For any $\left\{\phi_{m}\right\}\subset\mathcal{C}_{AB}$ and any probability distribution $\left\{p_{m}\right\}$, the convexity holds:
$C(\sum_{m}p_{m}\phi_{m})=\sum_{j}q_{j}C_{j}(\sum_{m}p_{m}\phi_{m})\leqslant\sum_{j}q_{j}\sum_{m}p_{m}C_{j}(\phi_{m})=\sum_{m}p_{m}\sum_{j}q_{j}C_{j}(\phi_{m})=\sum_{m}p_{m}C(\phi_{m}).$
Therefore, $C$ is an imaginarity measure. ∎
Property 3.
If $C$ is an imaginarity measure for channel $\phi$, then, quantifier $\hat{C}$ as:
$\hat{C}(\phi)\triangleq\min_{\left\{q_{j},\phi_{j}\right\}}\sum_{j}q_{j}C(\phi_{j}),$
is a proper measure for $\phi$, where $\min_{\left\{q_{j},\phi_{j}\right\}}$ represents the min of all possible channels decompositions as $\phi=\sum_{j}q_{j}\phi_{j}$ with $q_{j}\geqslant 0$.
Proof.
It can be seen that $\hat{C}(\phi)\geqslant 0$. When $\hat{C}(\phi)=0$, we can take the best decomposition of $\phi$, expressed as $\phi=\sum_{k}q_{k}\phi_{k}$, so that $\hat{C}(\phi)=\sum_{k}q_{k}C(\phi_{k})$. We can get $C(\phi_{k})=0$ since $q\geqslant 0$, that is, $\phi_{k}\in\mathcal{RC}$, thus $\phi\in\mathcal{RC}$, the faithfulness holds.
We can notice that, for any $\left\{q_{j}\right\}$, let every $\phi_{j}=\phi$, we can get $\hat{C}(\phi)=\min_{\left\{q_{j},\phi_{j}\right\}}\sum_{j}q_{j}C(\phi_{j})\leqslant C(\phi)$. For any quantum channel $\phi$, we take the best decomposition of $\phi$, expressed as $\phi=\sum_{k}q_{k}\phi_{k}$, so that $\hat{C}(\phi)=\sum_{k}q_{k}C(\phi_{k})$. For $\Theta\in\mathcal{RSC}$ with $\Theta(\cdot)=\sum_{m}M_{m}\cdot M_{m}^{\dagger}$, we denote $a_{k,m}=\text{Tr}(M_{m}\chi_{\phi_{k}}M_{m}^{\dagger})$, $p_{m}=\text{Tr}(M_{m}\chi_{\phi}M_{m}^{\dagger})$, then
$\hat{C}(\phi)=\sum_{k}q_{k}C(\phi_{k})$$\geqslant$ $\sum_{k}q_{k}\sum_{m}a_{k,m}C(\frac{M_{m}\chi_{\phi_{k}}M_{m}^{\dagger}}{a_{k,m}})=\sum_{m}p_{m}\sum_{k}\frac{q_{k}a_{k,m}}{p_{m}}C(\frac{M_{m}\chi_{\phi_{k}}M_{m}^{\dagger}}{a_{k,m}})$
$\geqslant$ $\sum_{m}p_{m}\sum_{k}C(\frac{q_{k}a_{k,m}}{p_{m}}\frac{M_{m}\chi_{\phi_{k}}M_{m}^{\dagger}}{a_{k,m}})=\sum_{m}p_{m}C(\frac{\sum_{k}q_{k}M_{m}\chi_{\phi_{k}}M_{m}^{\dagger}}{p_{m}})$
$=$ $\sum_{m}p_{m}C(\frac{M_{m}\chi_{\phi}M_{m}^{\dagger}}{p_{m}})\geqslant\sum_{m}p_{m}\hat{C}(\frac{M_{m}\chi_{\phi}M_{m}^{\dagger}}{p_{m}}).$
The first inequality is due to the strong monotonicity of $C$, the second inequality is due to the convexity of $C$ and $\sum_{k}\frac{q_{k}a_{k,m}}{p_{m}}=\frac{\sum_{k}q_{k}\text{Tr}(M_{m}\chi_{\phi_{k}}M_{m}^{\dagger})}{\text{Tr}(M_{m}\chi_{\phi}M_{m}^{\dagger})}=1$.
To verify convexity of $\hat{C}$ for simplicity, we only consider two quantum channels $\phi_{1}$ and $\phi_{2}$, and define $\psi=a\phi_{1}+b\phi_{2}$ with $a+b=1$. We take the best decomposition of $\phi_{1}$ and $\phi_{2}$ respectively as $\phi_{1}=\sum_{m}q_{1,m}\phi_{1,m}$ and $\phi_{2}=\sum_{n}q_{2,n}\phi_{2,n}$ so that the equations hold:
$\begin{cases}\hat{C}(\phi_{1})=\sum_{m}q_{1,m}C(\phi_{1,m}),\\ \hat{C}(\phi_{2})=\sum_{n}q_{2,n}C(\phi_{2,n}).\end{cases}$
Then, $a\hat{C}(\phi_{1})+b\hat{C}(\phi_{2})=\sum_{m}aq_{1,m}C(\phi_{1,m})+\sum_{n}bq_{2,n}C(\phi_{2,n})$.
For ease of writing, if $m\neq n$, we supplement $\left\{0,\bm{0}\right\}$ for the less set so that both sets have the same quantity, that is, $a\hat{C}(\phi_{1})+b\hat{C}(\phi_{2})=\sum_{m}aq_{1,m}C(\phi_{1,m})+bq_{2,m}C(\phi_{2,m})$. Substitute the new symbol $\left\{q_{2m-1},\psi_{2m-1}\right\}$ and $\left\{q_{2m},\psi_{2m}\right\}$ for $\left\{aq_{1,m},\phi_{1,m}\right\}$ and $\left\{bq_{2,m},\phi_{2,m}\right\}$ respectively. We have
$a\hat{C}(\phi_{1})+b\hat{C}(\phi_{2})=\sum_{m}q_{2m-1}C(\psi_{2m-1})+q_{2m}C(\psi_{2m})\geqslant\min_{q_{j},\psi_{j}}\sum_{j}q_{j}C(\psi_{j})=\hat{C}(\psi)=\hat{C}(a\phi_{1}+b\phi_{2}).$
In this way, it can be known that $\hat{C}$ is a proper imaginarity measure. ∎
3 Alternative framework for the imaginarity of quantum channels
In the previous section we refined the framework. However, the condition $\bm{(C_{3})}$ and $\bm{(C_{4})}$ are difficult to verify in some cases. Here, inspired by the similar alternative framework in many resource theory [22, 23], we have the following result.
Theorem 1.
$\bm{(C_{3})}+\bm{(C_{4})}$ is equivalent to $\bm{(C_{5})}$ when condition $\bm{(C_{1})}$ and $\bm{(C_{2})}$ holds. The condition $\bm{(C_{5})}$ is:
$\bm{(C_{5})}$. Additivity: For any $p_{1}$, $p_{2}$ satisfied $p_{1}+p_{2}=1$, any $\phi_{1}\in\mathcal{C}_{AB_{1}}$ and $\phi_{2}\in\mathcal{C}_{AB_{2}}$, let $\phi=p_{1}\phi_{1}\oplus p_{2}\phi_{2}$, $\phi\in\mathcal{C}_{AB}$ with $|B|=|B_{1}|+|B_{2}|$, then $C(\phi)=p_{1}C(\phi_{1})+p_{2}C(\phi_{2})$ holds.
Proof.
The proof is long and we divide it into steps.
$\bm{(C_{3})}+\bm{(C_{4})}\Rightarrow\bm{(C_{5})}$:
(STEP 1).
Let $Q_{1}=|0\rangle\langle 0|+\cdots+||B_{1}|-1\rangle\langle|B_{1}|-1|+||B|\rangle\langle|B||+\cdots+||B|+|B_{1}|-1\rangle\langle|B|+|B_{1}|-1|+|(|A|-1)|B|\rangle\langle(|A|-1)|B||+\cdots+|(|A|-1)|B|+|B_{1}|-1\rangle\langle(|A|-1)|B|+|B_{1}|-1|$,
$Q_{2}=||B_{1}|\rangle\langle|B_{1}||+\cdots+||B|-1\rangle\langle|B|-1|+||B|+|B_{1}|\rangle\langle|B|+|B_{1}||+\cdots+||B|+|B|-1\rangle\langle|B|+|B|-1|+|(|A|-1)|B|+|B_{1}|\rangle\langle(|A|-1)|B|+|B_{1}||+\cdots+|(|A|-1)|B|+|B|-1\rangle\langle(|A|-1)|B|+|B|-1|$.
We can give an intuitive expression.
$Q_{1}=\overbrace{\begin{pmatrix}I_{|B_{1}|}&&&&\\ &O_{|B_{2}|}&&&\\ &&\ddots&&\\ &&&I_{|B_{1}|}&\\ &&&&O_{|B_{2}|}\end{pmatrix}}^{|A||B|}=\bigoplus\limits_{j=0}^{|A|-1}\begin{pmatrix}I_{|B_{1}|}&\\ &O_{|B_{2}|}\end{pmatrix},$
$Q_{2}=\underbrace{\begin{pmatrix}O_{|B_{1}|}&&&&\\ &I_{|B_{2}|}&&&\\ &&\ddots&&\\ &&&O_{|B_{1}|}&\\ &&&&I_{|B_{2}|}\end{pmatrix}}_{|A||B|}=\bigoplus\limits_{j=0}^{|A|-1}\begin{pmatrix}O_{|B_{1}|}&\\ &I_{|B_{2}|}\end{pmatrix},$
where $I_{a}$ and $O_{b}$ denote identity matrix and zero matrix with dimensions $a$ and $b$ respectively.
Due to $Q_{1}^{\dagger}Q_{1}+Q_{2}^{\dagger}Q_{2}=I_{|A||B|}$ and $Q_{1}Q_{2}=O_{|A||B|}$, let $\Theta(\cdot)=Q_{1}\cdot Q_{1}^{\dagger}+Q_{2}\cdot Q_{2}^{\dagger}$, we can know that $\Theta\in\mathcal{RSC}_{ABA^{\prime}B^{\prime}}$.
Under the requirement of $\bm{(C_{5})}$, the Choi state of $\phi$ is
$\chi_{\phi}=\sum_{j,k}|j\rangle\langle k|\otimes\phi(|j\rangle\langle k|)/|A|=\sum_{j,k}|j\rangle\langle k|\otimes\big{[}p_{1}\phi_{1}(|j\rangle\langle k|)\oplus p_{2}\phi_{2}(|j\rangle\langle k|\big{]}/|A|.$
Thus,
$Q_{1}\chi_{\phi}Q_{1}^{\dagger}=\sum_{j,k}|j\rangle\langle k|\otimes\big{[}p_{1}\phi_{1}(|j\rangle\langle k|)\oplus\bm{0}_{2}(|j\rangle\langle k|\big{]}/|A|.$
It’s obvious that $\text{Tr}(Q_{1}\chi_{\phi}Q_{1}^{\dagger})$ is exactly $p_{1}$. Let $\chi_{\tilde{\phi}_{1}}=Q_{1}\chi_{\phi}Q_{1}^{\dagger}/p_{1}=\sum_{j,k}|j\rangle\langle k|\otimes\big{[}\phi_{1}(|j\rangle\langle k|)\oplus\bm{0}_{2}(|j\rangle\langle k|\big{]}/|A|$. Similarly we obtain $\chi_{\tilde{\phi}_{2}}$. At this point, the requirement of $\bm{(C_{3})}$ has been satisfied, that is
$\begin{cases}\Theta\in\mathcal{RSC}_{ABA^{\prime}B^{\prime}}\text{ with }\Theta(\cdot)=Q_{1}\cdot Q_{1}^{\dagger}+Q_{2}\cdot Q_{2}^{\dagger},\\ p_{m}=\text{Tr}(Q_{m}\chi_{\phi}Q_{m}^{\dagger}),\chi_{\tilde{\phi}_{m}}=Q_{m}\chi_{\phi}Q_{m}^{\dagger}/p_{m},\end{cases}$
we have
$C(\chi_{\phi})\geqslant p_{1}C(\chi_{\tilde{\phi}_{1}})+p_{2}C(\chi_{\tilde{\phi}_{2}}).$ (1)
(STEP 2).
For any $\rho\in\mathcal{D}_{A}$, $\rho=\sum_{j,k}\rho_{j,k}|j\rangle\langle k|$, any $\phi\in\mathcal{C}_{AB}$, $\phi(\rho)$ can be represented by $J_{\phi}$, in order to do this, $\rho$ and $J_{\phi}$ need to be tensor index realigned:
$\begin{cases}J^{\prime}_{\phi}=\sum_{j,k}\langle jk|\otimes\phi(|j\rangle\langle k|),\\ \rho^{\prime}=\sum_{j,k}\rho_{j,k}|jk\rangle,\end{cases}$
then $\phi(\rho)=J^{\prime}_{\phi}\rho^{\prime}=\sum_{j,k}\rho_{j,k}\phi(|j\rangle\langle k|)$. Under the requirement of $(C_{5})$, we have
$\phi(\rho)$ $=$ $\sum_{j,k}\rho_{j,k}\phi(|j\rangle\langle k|)=\sum_{j,k}\rho_{j,k}\big{[}p_{1}\phi_{1}(|j\rangle\langle k|)\oplus p_{2}\phi_{2}(|j\rangle\langle k|)\big{]}$
$=$ $p_{1}\sum_{j,k}\rho_{j,k}\big{[}\phi_{1}(|j\rangle\langle k|)\oplus\bm{0}_{2}(|j\rangle\langle k|)\big{]}+p_{2}\sum_{j,k}\rho_{j,k}\big{[}\bm{0}_{1}(|j\rangle\langle k|)\oplus\phi_{2}(|j\rangle\langle k|)\big{]}$
$=$ $p_{1}\tilde{\phi}_{1}(\rho)+p_{2}\tilde{\phi}_{2}(\rho),$
that is $\phi=p_{1}\tilde{\phi}_{1}+p_{2}\tilde{\phi}_{2}$. From $\bm{(C_{4})}$ We can know that
$C(\phi)\leqslant p_{1}C(\tilde{\phi}_{1})+p_{2}C(\tilde{\phi}_{2}).$ (2)
(STEP 3).
From (1) in (STEP 1) and (2) in (STEP 2), we obtain $C(\phi)=p_{1}C(\tilde{\phi}_{1})+p_{2}C(\tilde{\phi}_{2})$. Next we need to prove $C(\phi_{1})=C(\tilde{\phi}_{1})$, that is, we need to prove $C(\chi_{\phi_{1}})=C(\chi_{\tilde{\phi}_{1}})$.
First, we specify the range of index $i$ and $j$, let $i=0,1,\cdots,|B|-1$, $j=0,1,\cdots,|B_{1}|-1$. Let $K_{0}^{1}$ satisfies $\langle i|K_{0}^{1}|j\rangle=\delta_{i,j}$. An intuitive expression of $K_{0}^{1}$ is
$K_{0}^{1}=\left.\begin{pmatrix}I_{|B_{1}|}\\ \hat{O}\end{pmatrix}\right\}|B|,$
where $\hat{O}$ denotes zero matrix with dimension $|B_{2}|\times|B_{1}|$.
Let $K_{n}^{2}$ satisfy $\langle j|K_{n}^{2}|i\rangle=\delta_{i,j+n|B_{1}|}$, $n=0,1,\cdots,\left\lceil|B_{2}|/|B_{1}|\right\rceil$. $\left\lceil a\right\rceil$ is the ceil of $a$. We can give an example to make this easier to understand. See Example 1 after this proof.
Let
$\mathcal{K}_{0}^{1}=\bigoplus_{j=0}^{|A|-1}K_{0}^{1}=\begin{pmatrix}K_{0}^{1}&&\\ &\ddots&\\ &&K_{0}^{1}\end{pmatrix},\mathcal{K}_{n}^{2}=\bigoplus_{j=0}^{|A|-1}K_{n}^{2}=\begin{pmatrix}K_{n}^{2}&&\\ &\ddots&\\ &&K_{n}^{2}\end{pmatrix},$
we can verify $(\mathcal{K}_{0}^{1})^{\dagger}\mathcal{K}_{0}^{1}=I_{|B_{1}|}$ and $\sum_{j=0}^{\left\lceil|B_{2}|/|B_{1}|\right\rceil}(\mathcal{K}_{n}^{2})^{\dagger}\mathcal{K}_{n}^{2}=I_{|B|}$. Thus $\Theta_{1}\in\mathcal{RSC}_{AB_{1}AB}$ with $\Theta_{1}(\cdot)=\mathcal{K}_{0}^{1}\cdot(\mathcal{K}_{0}^{1})^{\dagger}$, $\Theta_{2}\in\mathcal{RSC}_{ABAB_{1}}$ with $\Theta_{2}(\cdot)=\sum_{j=0}^{\left\lceil|B_{2}|/|B_{1}|\right\rceil}\mathcal{K}_{n}^{2}\cdot(\mathcal{K}_{n}^{2})^{\dagger}$. We can also obtain $\Theta_{1}(\chi_{\phi_{1}})=\chi_{\tilde{\phi}_{1}}$ and $\Theta_{2}(\chi_{\tilde{\phi}_{1}})=\chi_{\phi_{1}}$.
Combining with $\bm{(C_{2})}$, the above implies that the following inequality holds:
$C(\chi_{\phi_{1}})=C(\Theta_{2}(\chi_{\tilde{\phi}_{1}}))\leqslant C(\chi_{\tilde{\phi}_{1}})=C(\Theta_{1}(\chi_{\phi_{1}}))\leqslant C(\chi_{\phi_{1}}).$
As a result, $C(\chi_{\phi_{1}})=C(\chi_{\tilde{\phi}_{1}})$. It follows from a similar discussion that $C(\chi_{\phi_{2}})=C(\chi_{\tilde{\phi}_{2}})$.
Up to now, we prove $\bm{(C_{5})}$ holds: $C(\phi)=p_{1}C(\phi_{1})+p_{2}C(\phi_{2})$.
$\bm{(C_{5})}\Rightarrow\bm{(C_{3})}+\bm{(C_{4})}$:
(STEP 4).
For $\Theta\in\mathcal{RSC}_{ABA^{\prime}B^{\prime}}$ with $\Theta(\cdot)=\sum_{m}M_{m}\cdot M_{m}^{\dagger}$, we introduce an auxiliary system $S$ with dimension $N$ and orthonormal basis $\left\{|n\rangle\right\}_{n=0}^{N-1}$. The system $AB$ formed by Choi state and the auxiliary system $S$ form a combined system $ABS$, of which the basis is $\left\{|j\alpha\rangle\otimes|n\rangle\right\}$ and the whole system is initially in the state $\chi_{\phi}\otimes|0\rangle\langle 0|\triangleq\chi_{\phi}^{ABS}$.
Let $U_{n}=\sum_{k=0}^{N-1}|(k+n)\text{ mod }N\rangle\langle k|$, then $\Theta^{ABS}(\cdot)\in\mathcal{RSC}_{ABSA^{\prime}B^{\prime}S}$ with $\Theta^{ABS}(\cdot)=\sum_{n}(M_{n}\otimes U_{n})\cdot(M_{n}\otimes U_{n})^{\dagger}$. Thus
$\Theta^{ABS}(\chi_{\phi}^{ABS})=\sum_{n}(M_{n}\otimes U_{n})(\chi_{\phi}\otimes|0\rangle\langle 0|)(M_{n}\otimes U_{n})^{\dagger}=\sum_{n}M_{n}\chi_{\phi}M_{n}^{\dagger}\otimes|n\rangle\langle n|\triangleq\sum_{n}p_{n}\chi_{\phi_{n}}\otimes|n\rangle\langle n|,$
where $p_{n}=\text{Tr}(M_{n}\chi_{\phi}M_{n}^{\dagger})$ and $\chi_{\phi_{n}}=M_{n}\chi_{\phi}M_{n}^{\dagger}/p_{n}$.
We can know that $\sum_{n}p_{n}=1$, $\phi_{n}\in\mathcal{C}_{AB_{n}}$ with $|B|=\sum_{n}|B_{n}|$. From $\bm{(C_{5})}$, we get
$\begin{cases}C(\chi_{\phi}^{ABS})=C(\chi_{\phi}\otimes|0\rangle\langle 0|)=C(\chi_{\phi}),\\ C(\Theta^{ABS}(\chi_{\phi}^{ABS}))=C(\sum_{n}p_{n}\chi_{\phi_{n}}\otimes|n\rangle\langle n|)=C(\bigoplus_{n}p_{n}\chi_{\phi_{n}})=\sum_{n}p_{n}C(\chi_{\phi_{n}}).\end{cases}$
From $\bm{(C_{2})}$, we can obtain
$C(\chi_{\phi})=C(\chi_{\phi}^{ABS})\geqslant C(\Theta^{ABS}(\chi_{\phi}^{ABS}))=\sum_{n}p_{n}C(\chi_{\phi_{n}}),$
that is, condition $\bm{(C_{3})}$ holds.
(STEP 5).
For any $\left\{\phi_{m}\right\}\subset\mathcal{C}_{AB}$ and any probability distribution $\left\{p_{m}\right\}$, we construct the similar auxiliary system $S$ with orthonormal basis $\left\{|m\rangle\right\}_{m}$ and define $\chi_{\phi}^{ABS}=\sum_{m}p_{m}\chi_{\phi_{m}}\otimes|m\rangle\langle m|$. Due to $\sum_{m}(I_{|A||B|}\otimes|0\rangle\langle m|)^{\dagger}(I_{|A||B|}\otimes|0\rangle\langle m|)=I_{|A||B||S|}$, $\Theta^{ABS}\in\mathcal{RSC}_{ABSA^{\prime}B^{\prime}S}$ with $\Theta^{ABS}(\cdot)=\sum_{m}(I_{|A||B|}\otimes|0\rangle\langle m|)\cdot(I_{|A||B|}\otimes|0\rangle\langle m|)^{\dagger}$. Then
$\Theta^{ABS}(\chi_{\phi}^{ABS})=\sum_{m}(I_{|A||B|}\otimes|0\rangle\langle m|)\chi_{\phi}^{ABS}(I_{|A||B|}\otimes|0\rangle\langle m|)^{\dagger}=\sum_{m}p_{m}\chi_{\phi_{m}}\otimes|0\rangle\langle 0|.$
We know from $\bm{(C_{2})}$ and $\bm{(C_{5})}$ that the following inequality holds:
$\sum_{m}p_{m}C(\chi_{\phi_{m}})=C(\chi_{\phi}^{ABS})\geqslant C(\Theta^{ABS}(\chi_{\phi}^{ABS}))=C(\sum_{m}p_{m}\chi_{\phi_{m}}),$
that is, condition $\bm{(C_{4})}$ holds.
Therefore, the proof is finished. ∎
Example 1.
Suppose $|B_{1}|=2$ and $|B_{2}|=3$, then $i=0,1,\cdots,4$, $j=0,1$, $n=0,1,2$ and
$K_{0}^{2}=\begin{pmatrix}1&0&0&0&0\\ 0&1&0&0&0\end{pmatrix}\triangleq\begin{pmatrix}I_{|B_{1}|}&O_{|B_{1}|}&\tilde{O}\end{pmatrix},$
$K_{1}^{2}=\begin{pmatrix}0&0&1&0&0\\ 0&0&0&1&0\end{pmatrix}\triangleq\begin{pmatrix}O_{|B_{1}|}&I_{|B_{1}|}&\tilde{O}\end{pmatrix},$
$K_{2}^{2}=\begin{pmatrix}0&0&0&0&1\\ 0&0&0&0&0\end{pmatrix}\triangleq\begin{pmatrix}O_{|B_{1}|}&O_{|B_{1}|}&\tilde{I}\end{pmatrix},$
where $\tilde{O}$ denotes the first $(|B_{2}|\text{ mod }|B_{1}|)$ columns of $O_{|B_{1}|}$, same does $\tilde{I}$.
It is often difficult to verify whether a measure has strong monotonicity and convexity directly, which may be easier after using Theorem 1. When verifying whether a quantifier is a proper measure, we can choose verifying $\bm{(C_{1})}$$\bm{(C_{4})}$ or verifying $\bm{(C_{1})}+\bm{(C_{2})}+\bm{(C_{5})}$.
Under the requirement of $\bm{(C_{5})}$, that is, for any $p_{1}$, $p_{2}$ satisfied $p_{1}+p_{2}=1$, any $\phi_{1}\in\mathcal{C}_{AB_{1}}$ and $\phi_{2}\in\mathcal{C}_{AB_{2}}$, let $\phi=p_{1}\phi_{1}\oplus p_{2}\phi_{2}$, $\phi\in\mathcal{C}_{AB}$ with $|B|=|B_{1}|+|B_{2}|$. We have $\chi_{\phi_{1}}$ and $\chi_{\phi_{2}}$, and we define $\chi_{\tilde{\phi}_{1}}$ and $\chi_{\tilde{\phi}_{2}}$ below.
$\begin{cases}\chi_{\phi_{1}}=\sum_{j,k}|j\rangle\langle k|\otimes\phi_{1}(|j\rangle\langle k|)/|A|,\\ \chi_{\phi_{2}}=\sum_{j,k}|j\rangle\langle k|\otimes\phi_{2}(|j\rangle\langle k|)/|A|,\\ \chi_{\tilde{\phi}_{1}}=\sum_{j,k}|j\rangle\langle k|\otimes\big{[}\phi_{1}(|j\rangle\langle k|)\oplus\bm{0}_{2}(|j\rangle\langle k|\big{]}/|A|,\\ \chi_{\tilde{\phi}_{2}}=\sum_{j,k}|j\rangle\langle k|\otimes\big{[}\bm{0}_{1}(|j\rangle\langle k|)\oplus\phi_{2}(|j\rangle\langle k|\big{]}/|A|.\end{cases}$
Thus the following property holds.
Property 4.
$C$ is a proper imaginarity measure for quantum channels, the equation holds:
$C(p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}})=C(p_{1}\chi_{\tilde{\phi}_{1}}+p_{2}\chi_{\tilde{\phi}_{2}})=C(\chi_{\phi}).$ (3)
The proof of Property 4 is obvious, here is a helpful example.
Example 2.
The following matrix transformations hold and they are invertible.
$\begin{pmatrix}a_{11}&a_{12}&&&&\\ a_{21}&a_{22}&&&&\\ &&a_{33}&&&\\ &&&b_{11}&b_{12}&\\ &&&b_{21}&b_{22}&\\ &&&&&b_{33}\end{pmatrix}\xrightarrow{K_{1}(\cdot)K_{1}^{\dagger}}\begin{pmatrix}a_{11}&&a_{12}&&&\\ &a_{33}&&&&\\ a_{21}&&a_{22}&&&\\ &&&b_{22}&b_{21}&\\ &&&b_{12}&b_{11}&\\ &&&&&b_{33}\end{pmatrix}\xrightarrow{K_{2}(\cdot)K_{2}^{\dagger}}\begin{pmatrix}a_{11}&&a_{12}&&&\\ &b_{11}&&b_{12}&&\\ a_{21}&&a_{22}&&&\\ &b_{21}&&b_{22}&&\\ &&&&a_{33}&\\ &&&&&b_{33}\\ \end{pmatrix},$
where
$K_{1}=\begin{pmatrix}1&&&&&\\ &&1&&&\\ &1&&&&\\ &&&&1&\\ &&&1&&\\ &&&&&1\\ \end{pmatrix},K_{2}=\begin{pmatrix}1&&&&&\\ &&&&1&\\ &&1&&&\\ &&&1&&\\ &1&&&&\\ &&&&&1\\ \end{pmatrix}.$
Due to $K^{\dagger}_{j}K_{j}=I$, then the invertible $\Theta_{j}\in\mathcal{RSC}$ with $\Theta_{j}(\cdot)=K_{j}\cdot K_{j}^{\dagger}$, $j=1,2$.
Just like $\Theta_{1}$ and $\Theta_{2}$ in this example, we can construct multiple invertible channels such that
$\Theta_{1}(\Theta_{2}(\cdots\Theta_{k}(p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}})\cdots))=p_{1}\chi_{\tilde{\phi}_{1}}+p_{2}\chi_{\tilde{\phi}_{2}}.$
In this way, we can know that the equation (3) holds from the condition $\bm{(C_{2})}$.
4 Imaginarity measures of quantum channels
The Choi theorem [27] tell us that the Choi matrix of the channel $\phi$ is positive semidefinite, so that $\chi_{\phi}\geqslant 0$, we can see that, for any quantum channel $\phi$ with $\phi(\cdot)=\sum_{n}K_{n}\cdot K_{n}^{\dagger}$, its Choi state $\chi_{\phi}=\sum_{j,k}|j\rangle\langle k|\otimes\phi(|j\rangle\langle k|)/|A|$, thus the trace:
$\text{Tr}(\chi_{\phi})=\text{Tr}\Big{[}\sum_{j,k}|j\rangle\langle k|\otimes\phi(|j\rangle\langle k|)/|A|\Big{]}=\frac{1}{|A|}\sum_{j}\text{Tr}\Big{[}\phi(|j\rangle\langle j|)\Big{]}=\frac{1}{|A|}\sum_{j}1=1.$
In this way, we can embed Choi states into quantum states:
$\mathcal{C}_{AB}\hookrightarrow\mathcal{D}_{AB}.$
It seems then that many imaginarity measures for quantum states can be equally applied to quantum channels. Thus in this section, we introduce three imaginarity measures of quantum channels, that is, measures via on the robustness, the trace norm and the sandwiched Rényi relative entropy, respectively. We can exactly verify these three measures in two ways.
4.1 Measure via on the robustness
The Measure via on the robustness is appeared in many resource theories [1, 18]. Similar to these, we define the imaginarity measure of quantum channels as follow.
$C_{r}(\phi)=\min_{\tilde{\phi}\in\mathcal{RC}_{AB}}\left\{s\geqslant 0:\frac{\chi_{\phi}+s\chi_{\tilde{\phi}}}{1+s}\triangleq\chi_{\psi},\psi\in\mathcal{RC}_{AB}\right\}$
We have the following result.
Theorem 2.
$C_{r}(\phi)$ is a proper imaginarity measure of quantum channels.
Before our proof, we need to introduce pseudomixture [28, 29, 30, 7] of $\chi_{\phi}$, that is,
$\chi_{\phi}=[1+C_{r}(\chi_{\phi})]\chi_{\phi_{1}}-C_{r}(\chi_{\phi})\chi_{\phi_{2}},$
where $\phi_{1}\in\mathcal{RC}_{AB}$, and $\phi_{2}$ is the optimal channel of the $\min$ in the definition of $C_{r}(\phi)$.
Proof.
We only verify $\bm{(C_{1})}$+$\bm{(C_{3})}$+$\bm{(C_{4})}$ since $\bm{(C_{2})}$ can be derived.
$\bm{(C_{1})}$:
Obviously $C_{r}(\phi)\geqslant 0$. If $C_{r}(\phi)=0$, that means $\frac{\chi_{\phi}+0}{1+0}=\chi_{\psi}$, that is $\phi=\psi\in\mathcal{RC}_{AB}$. The converse is certainly true.
$\bm{(C_{3})}$:
Let $\Theta\in\mathcal{RSC}_{ABA^{\prime}B^{\prime}}$ with $\Theta(\cdot)=\sum_{m}M_{m}\cdot M_{m}^{\dagger}$, for any quantum channel $\phi$ with its pseudomixture, under the requirement of $\bm{(C_{3})}$, $p_{m}=\text{Tr}(M_{m}\chi_{\phi}M_{m}^{\dagger})$ and $\chi_{\phi_{m}}=M_{m}\chi_{\phi}M_{m}^{\dagger}/p_{m}$. Then we obtain
$M_{m}\chi_{\phi}M_{m}^{\dagger}=[1+C_{r}(\chi_{\phi})]M_{m}\chi_{\phi_{1}}M_{m}^{\dagger}-C_{r}(\chi_{\phi})M_{m}\chi_{\phi_{2}}M_{m}^{\dagger}.$
Let
$\begin{cases}\chi_{\phi_{1}}^{m}=\frac{1}{1+s_{m}}\frac{1}{p_{m}}[1+C_{r}(\chi_{\phi})]M_{m}\chi_{\phi_{1}}M_{m}^{\dagger},\\ \chi_{\phi_{2}}^{m}=\frac{1}{s_{m}}\frac{1}{p_{m}}C_{r}(\chi_{\phi})M_{m}\chi_{\phi_{2}}M_{m}^{\dagger},\end{cases}$
where $s_{m}=\frac{1}{p_{m}}C(\chi_{\phi})\text{Tr}(M_{m}\chi_{\phi_{1}}M_{m}^{\dagger})$.
We get $\chi_{\phi_{m}}=(1+s_{m})\chi_{\phi_{1}}^{m}-s_{m}\chi_{\phi_{2}}^{m}$, that is, $\frac{\chi_{\phi_{m}}+s_{m}\chi_{\phi_{2}}^{m}}{1+s_{m}}=\chi_{\phi_{1}}^{m}$. Thus we get $C_{r}(\chi_{\phi_{m}})\leqslant s_{m}$ from definition, that is, $p_{m}C_{r}(\chi_{\phi_{m}})\leqslant C(\chi_{\phi})\text{Tr}(M_{m}\chi_{\phi_{1}}M_{m}^{\dagger})$.
Therefore, we obtain
$\sum_{m}p_{m}C_{r}(\chi_{\phi_{m}})$$\leqslant$ $\sum_{m}C(\chi_{\phi})\text{Tr}(M_{m}\chi_{\phi_{1}}M_{m}^{\dagger})$
$=$ $C(\chi_{\phi})\sum_{m}\text{Tr}(M_{m}^{\dagger}M_{m}\chi_{\phi_{1}})$
$=$ $C(\chi_{\phi})\text{Tr}(\sum_{m}M_{m}^{\dagger}M_{m}\chi_{\phi_{1}})$
$=$ $C(\chi_{\phi})$
This means the condition $\bm{(C_{3})}$ holds.
$\bm{(C_{4})}$:
For simplicity, we only consider two quantum channels $\phi_{1}$ and $\phi_{2}$, and define $\phi=p\phi_{1}+(1-p)\phi_{2}$ with any $p\in(0,1)$, that is, $\chi_{\phi}=p\chi_{\phi_{1}}+(1-p)\chi_{\phi_{2}}$. The measurements $C_{r}$ for $\phi_{1}$ and $\phi_{2}$ are $C_{r}(\chi_{\phi_{1}})$ and $C_{r}(\chi_{\phi_{2}})$. Let $s_{p}=pC_{r}(\chi_{\phi_{1}})+(1-p)C_{r}(\chi_{\phi_{2}})$.
Consider the pseudomixtures of $\chi_{\phi_{1}}$ and $\chi_{\phi_{2}}$:
$\chi_{\phi_{1}}=[1+C_{r}(\chi_{\phi_{1}})]\chi_{\phi_{1,1}}-C_{r}(\chi_{\phi_{1}})\chi_{\phi_{1,2}}\mbox{ and }\chi_{\phi_{2}}=[1+C_{r}(\chi_{\phi_{2}})]\chi_{\phi_{2,1}}-C_{r}(\chi_{\phi_{2}})\chi_{\phi_{2,2}},$
then let
$\begin{cases}\chi_{\psi}=\frac{p[1+C_{r}(\chi_{\phi_{1}})]\chi_{\phi_{1,1}}+(1-p)[1+C_{r}(\chi_{\phi_{2}})]\chi_{\phi_{2,1}}}{1+s_{p}},\\ \chi_{\tilde{\phi}}=\frac{pC_{r}(\chi_{\phi_{1}})\chi_{\phi_{1,2}}+(1-p)C_{r}(\chi_{\phi_{2}})\chi_{\phi_{2,2}}}{s_{p}}.\end{cases}$
We have $\chi_{\phi}=(1+s_{p})\chi_{\psi}-s_{p}\chi_{\tilde{\phi}}$, that is, $\chi_{\psi}=\frac{\chi_{\phi}+s_{p}\chi_{\tilde{\phi}}}{1+s_{p}}$, thus we get $s_{p}\geqslant C_{r}(\chi_{\phi})$.
Therefore, $pC_{r}(\chi_{\phi_{1}})+(1-p)C_{r}(\chi_{\phi_{2}})=s_{p}\geqslant C_{r}(\chi_{\phi})=C(p\chi_{\phi_{1}}+(1-p)\chi_{\phi_{2}})$. The condition $\bm{(C_{4})}$ holds. ∎
4.2 Measure via on the trace norm
To obtain the channel is real, we can consider the trace norm:
$C_{t}(\phi)=||\chi_{\phi}-\chi_{\phi}^{*}||_{\text{Tr}}=\text{Tr}|\chi_{\phi}-\chi_{\phi}^{*}|.$
where $\chi_{\phi}^{*}$ means the conjugate of $\chi_{\phi}$ and $|\chi_{\phi}-\chi_{\phi}^{*}|=\sqrt{(\chi_{\phi}-\chi_{\phi}^{*})^{\dagger}(\chi_{\phi}-\chi_{\phi}^{*})}$.
Theorem 3.
$C_{t}(\phi)$ is a proper imaginarity measure of quantum channels.
Proof.
The condition $\bm{(C_{1})}$ holds due to the definition of the trace norm.
There is $||\chi_{\phi}-\chi_{\phi}^{*}||_{\text{Tr}}\geqslant||\Theta(\chi_{\phi})-\Theta(\chi_{\phi}^{*})||_{\text{Tr}}$ for any CPTP map [31], and it certainly holds for real superchannels. Since $\Theta$ is a real superchannel, we can know that $[\Theta(\chi_{\phi})]^{*}=\Theta(\chi_{\phi}^{*})$ due to
$[\Theta(\chi_{\phi})]^{*}=\left[\sum_{m}M_{m}\chi_{\phi}M_{n}^{\dagger}\right]^{*}=\sum_{m}M_{m}\chi_{\phi}^{*}M_{n}^{\dagger}=\Theta(\chi_{\phi}^{*}).$
Thus $\bm{(C_{2})}$ is satisfied.
Under the requirement of $\bm{(C_{5})}$, from Property 4, we have:
$C_{t}(\chi_{\phi})$ $=$ $C_{t}(p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}})=||(p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}})-(p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}})^{*}||_{\text{Tr}}$
$=$ $||(p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}})-(p_{1}\chi_{\phi_{1}}^{*}\oplus p_{2}\chi_{\phi_{2}}^{*})||_{\text{Tr}}=||p_{1}(\chi_{\phi_{1}}-\chi_{\phi_{1}}^{*})\oplus p_{2}(\chi_{\phi_{2}}-\chi_{\phi_{2}}^{*})||_{\text{Tr}}$
$=$ $p_{1}||(\chi_{\phi_{1}}-\chi_{\phi_{1}}^{*})||_{\text{Tr}}+p_{2}||(\chi_{\phi_{2}}-\chi_{\phi_{2}}^{*})||_{\text{Tr}}=p_{1}C_{t}(\chi_{\phi_{1}})+p_{2}C_{t}(\chi_{\phi_{2}}).$
Therefore, $C_{r}(\phi)$ satisfies $\bm{(C_{1})}$+$\bm{(C_{2})}$+$\bm{(C_{5})}$. ∎
4.3 Measure via on the sandwiched Rényi relative entropy
The sandwiched Rényi relative entropy [32] defined for quantum states with the parameter $\alpha$ as follow:
$D_{\alpha}(\rho||\sigma)=\frac{1}{\alpha-1}\text{log }\text{Tr}(\sigma^{\frac{1-\alpha}{2\alpha}}\rho\sigma^{\frac{1-\alpha}{2\alpha}})^{\alpha},$
where Now, we will introduce the imaginarity measure of quantum channels via on the sandwiched Rényi relative entropy, that is,
$C_{\alpha}(\phi)=1-\text{Tr}\left[(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\chi_{\phi}(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\right]^{\alpha},$
where $\frac{1}{2}\leqslant\alpha<1$. We have the following results.
Theorem 4.
$C_{\alpha}(\phi)$ is a proper imaginarity measure of quantum channels.
Proof.
It is shown [32] that $D_{\alpha}(\rho||\sigma)\geqslant 0$ when $\frac{1}{2}\leqslant\alpha<1$ and equation holds if and only if $\rho=\sigma$, that is, $\text{Tr}(\sigma^{\frac{1-\alpha}{2\alpha}}\rho\sigma^{\frac{1-\alpha}{2\alpha}})^{\alpha}\leqslant 1$. Thus we consider Choi state as a quantum state and then we have, $C_{\alpha}(\phi)\geqslant 0$. The equation holds if and only if $\chi_{\phi}=\chi_{\phi}^{*}$. The condition $\bm{(C_{1})}$ is satisfied.
From [32], we have:
$D_{\alpha}(\chi_{\phi}||\chi_{\phi}^{*})\geqslant D_{\alpha}\left[\Theta(\chi_{\phi})||\Theta(\chi_{\phi}^{*})\right]$
$\iff$ log $\text{log }\text{Tr}\left[(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\chi_{\phi}(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\right]^{\alpha}\leqslant\text{log }\text{Tr}\left\{[\Theta(\chi_{\phi}^{*})]^{\frac{1-\alpha}{2\alpha}}\Theta(\chi_{\phi})[\Theta(\chi_{\phi}^{*})]^{\frac{1-\alpha}{2\alpha}}\right\}^{\alpha}$
$\iff$ $\text{Tr}\left[(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\chi_{\phi}(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\right]^{\alpha}\leqslant\text{Tr}\left\{[\Theta(\chi_{\phi}^{*})]^{\frac{1-\alpha}{2\alpha}}\Theta(\chi_{\phi})[\Theta(\chi_{\phi}^{*})]^{\frac{1-\alpha}{2\alpha}}\right\}^{\alpha}$
$\iff$ $\text{Tr}\left[(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\chi_{\phi}(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\right]^{\alpha}\leqslant\text{Tr}\left\{([\Theta(\chi_{\phi})]^{*})^{\frac{1-\alpha}{2\alpha}}\Theta(\chi_{\phi})([\Theta(\chi_{\phi})]^{*})^{\frac{1-\alpha}{2\alpha}}\right\}^{\alpha}$
$\iff$ $C_{\alpha}(\phi)\geqslant C_{\alpha}\left[\Theta(\phi)\right].$
Thus $\bm{(C_{2})}$ holds.
Under the requirement of $\bm{(C_{5})}$, we have:
$\text{Tr}\left[(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\chi_{\phi}(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\right]^{\alpha}$ $=$ $\text{Tr}\left\{\left[(p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}})^{*}\right]^{\frac{1-\alpha}{2\alpha}}(p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}})\left[(p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}})^{*}\right]^{\frac{1-\alpha}{2\alpha}}\right\}^{\alpha}$
$=$ $\text{Tr}\left\{\bigoplus_{j=1}^{2}p_{j}^{\frac{1-\alpha}{2\alpha}}p_{j}p_{j}^{\frac{1-\alpha}{2\alpha}}(\chi_{\phi_{j}}^{*})^{\frac{1-\alpha}{2\alpha}}\chi_{\phi_{j}}(\chi_{\phi_{j}}^{*})^{\frac{1-\alpha}{2\alpha}}\right\}^{\alpha}$
$=$ $\sum_{j=1}^{2}p_{j}\text{Tr}\left[(\chi_{\phi_{j}}^{*})^{\frac{1-\alpha}{2\alpha}}\chi_{\phi_{j}}(\chi_{\phi_{j}}^{*})^{\frac{1-\alpha}{2\alpha}}\right]^{\alpha}$
In this way, we obtain:
$C_{\alpha}(\phi)=1-\text{Tr}\left[(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\chi_{\phi}(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\right]^{\alpha}=\sum_{j=1}^{2}p_{j}\left\{1-\text{Tr}\left[(\chi_{\phi_{j}}^{*})^{\frac{1-\alpha}{2\alpha}}\chi_{\phi_{j}}(\chi_{\phi_{j}}^{*})^{\frac{1-\alpha}{2\alpha}}\right]^{\alpha}\right\}=p_{1}C_{\alpha}(\phi_{1})+p_{2}C_{\alpha}(\phi_{2})$
The condition $\bm{(C_{5})}$ holds. ∎
5 Conclusion
We refine the framework of imaginarity of quantum channels which originally only has $\bm{(C_{1})}$ and $\bm{(C_{2})}$ for its measure. We add two conditions, strong monotonicity $\bm{(C_{3})}$ and convexity $\bm{(C_{4})}$. Besides, we give some properties of imaginarity measure of quantum channels. Sometimes it is not easy to verify strong monotonicity and convexity, so we introduce an alternative framework, that is, $\bm{(C_{3})}+\bm{(C_{4})}$ is equivalent to $\bm{(C_{5})}$ when $\bm{(C_{1})}$ and $\bm{(C_{2})}$ hold. This is analogous to imaginarity or coherence of quantum states. Since Choi states can embed quantum states, we also think that imaginarity measures of quantum states can also be used for quantum channels and present three imaginarity measures for quantum channels, they are $C_{r}$, $C_{t}$ and $C_{\alpha}$ via on the robustness, the trace norm and the sandwiched Rényi relative entropy, respectively.
Acknowledgement
We thank Jianwei Xu for helpful discussions and comments. This work was supported by National Natural Science Foundation of China (Grants No. 12271474).

### v2

@@
 Abstract
-At present, quantum channels have been widely concerned, and many ways to quantify quantum channels have been proposed, which has led to the generation of many resource theories for quantum channels.
-We add strong monotonicity and convexity to the requirement of imaginarity measure of quantum channels to make the measure proper.
-We also introduce an alternative framework to simplify the process of verifying whether a quantifier is a proper measure.
-We present three imaginarity measures of quantum channels via on the robustness, the trace norm and entropy, respectively.
+In this paper, we introduce the framework for quantifying the imaginarity of quantum channels.
+Besides, an alternative framework is given together to simplify the process of verifying the condition.
+We present three imaginarity measures of quantum channels based on the robustness, the trace norm, and the entropy, respectively.
 Some properties are also given.
-Keywords: quantum resource theories, quantum channels, imaginarity measures, alternative framework
+keywords:
+quantum resource theories , quantum channels , imaginarity measures , alternative framework
+[…]
+organization=School of Mathematics, addressline=Harbin Institute of Technology, city=Harbin, postcode=150001, country=China
 1 Introduction
-Imaginary numbers provide a broad perspective and additional methods for modern scientific research.
-Quantum resource theory can establish new connections among various research directions.
-With the development of quantum resource theory, the imaginarity resource theory [1] proposed in recent years has attracted increasing attention.
-Some examples of applications, the imaginarity plays a crucial role in state discrimination [2], the imaginarity operation is more advantageous to distinguish the quantum channel without the help of the auxiliary system [3].
-Many imaginarity measures have been found for quantum states, the fidelity [2], the convex roof [4], the geometric [3] and some imaginarity measures induced by entropies [5, 6], some of which can establish relations due to their own mathematical properties [6, 7].
-Just like quantum states, quantum channels also deserve attention.
-Some studies begin to investigate the properties of quantum channels [8, 9, 10, 11, 12].
-Resource theories of quantum channels was established in [13], and many resource theories of quantum channels have been proposed, such as coherence of quantum channels [14], entanglement of quantum channels [15] and so on [16, 17].
-Based on this, more measures for quantum channels have been found [18, 19, 20, 21].
-In order to simplify the process of verifying whether a quantifier can be a coherence measure, alternative framework for quantifying coherence have been proposed [22] in the coherence resource theory.
-Similar alternative frameworks can be established in many other resource theories, such as imaginarity resource theory [7], coherence of quantum channels [23].
-The concept of the imaginarity of quantum channels was first introduced in [24], and the definitions of free channels and free superchannels have been visited again recently [25].
-It can be seen that the concept is worthing considing.
-Similar to the coherence of quantum channels, in this paper, we refine the imaginarity of quantum channels and introduce alternative framework to facilitate the verification of whether a quantifier can be a suitable measure.
-We also show three imaginarity measures of quantum channels.
-Some properties of the measure are discussed.
-2 Refinement for the imaginarity of quantum channels
+Imaginary numbers offer a broad perspective and provide additional methodologies for modern scientific research.
+Quantum resource theory facilitates the establishment of new connections across various research directions.
+In recent years, the concept of imaginarity resource theory [1], which has emerged alongside the development of quantum resource theory, has garnered increasing attention.
+Several applications can illustrate its significance.
+For instance, imaginarity plays a critical role in state discrimination [2].
+The imaginarity operation proves to be particularly advantageous for distinguishing quantum channels without the need for an auxiliary system [3].
+Numerous imaginarity measures have been identified for quantum states, including those based on fidelity [2], the convex roof [4], geometric considerations [3], and some induced by entropies [5, 6].
+Some of them establish relations due to their mathematical properties [6, 7].
+Just as quantum states warrant attention, quantum channels also merit investigation.
+Recent studies have begun to explore the properties of quantum channels [8, 9, 10, 11, 12].
+The resource theories of quantum channels were first established in [13], with several subsequent theories proposed, such as coherence in quantum channels [14], entanglement in quantum channels [15], and more [16, 17].
+Building upon these frameworks, additional measures for quantum channels have been introduced [18, 19, 20, 21].
+The concept of the imaginarity of quantum channels was first introduced in [24], and the definitions of free channels and free superchannels have been revisited recently [25].
+This indicates that the concept is of significant interest and worth considering.
+The structure of this paper is as follows: Similar to the coherence of quantum channels, we refine the concept of imaginarity for quantum channels and present an alternative framework that facilitates the verification of whether a given quantifier can serve as a suitable measure.
+We discuss the properties of imaginarity measures of quantum channels.
+Additionally, we propose three measures of the imaginarity of quantum channels based on the robustness, the trace norm, and the entropy, respectively.
+2 Framework for quantifying the imaginarity of quantum channels
 Let $H_{A}$ and $H_{B}$ be two Hilbert space with dimensions $|A|$ and $|B|$, and orthonormal basis $\left\{|j\rangle\right\}_{j=0}^{|A|-1}=\left\{|k\rangle\right\}_{k=0}^{|A|-1}$ and $\left\{|\alpha\rangle\right\}_{\alpha=0}^{|B|-1}=\left\{|\beta\rangle\right\}_{\beta=0}^{|B|-1}$.
@@
 A quantum channel $\phi\in\mathcal{C}_{AB}$ can be represented by Choi matrix,
-$J_{\phi}=\sum_{j,k}|j\rangle\langle k|\otimes\phi(|j\rangle\langle k|)=\sum_{j,k,\alpha,\beta}\phi_{j,k,\alpha,\beta}|j\rangle\langle k|\otimes|\alpha\rangle\langle\beta|$
+$J_{\phi}$ $=$ $\sum_{j,k}|j\rangle\langle k|\otimes\phi(|j\rangle\langle k|)$
+$=$ $\sum_{j,k,\alpha,\beta}\phi_{j,k,\alpha,\beta}|j\rangle\langle k|\otimes|\alpha\rangle\langle\beta|$
 with $\phi_{j,k,\alpha,\beta}=\langle\alpha|\phi(|j\rangle\langle k|)|\beta\rangle$.
 The Choi state of $\phi$ is $\frac{J_{\phi}}{|A|}\triangleq\chi_{\phi}$.
-We call channel $\phi$ a real channel if $\chi_{\phi}$ is an real matrix, and denote $\mathcal{RC}_{AB}$ the set of all the real channels.
+We call channel $\phi$ a real channel if $\chi_{\phi}$ is a real matrix, and denote $\mathcal{RC}_{AB}$ the set of all the real channels.
 The free channel is defined as the real channel.
-This is exactlt the free operation in imaginarity resource theory, thus it is appropriate to define the free channel by Choi state.
-Let $\mathcal{SC}_{ABA^{\prime}B^{\prime}}$ be the sets of all superchannels from $\mathcal{C}_{AB}$ to $\mathcal{C}_{A^{\prime}B^{\prime}}$.
-It can be known that [14, 26] superchannel $\Theta$ also has the expression of Kraus operators $\left\{M_{m}\right\}_{m}$ with $\chi_{\Theta(\phi)}=\sum_{m}M_{m}\chi_{\phi}M_{m}^{\dagger}\triangleq\tilde{\Theta}(\chi_{\phi})$, where $\tilde{\Theta}$ is the channel corresponding to the superchannel.
-We will not distinguish the signs, $\Theta$ or $\tilde{\Theta}$, since it can be judged by the context.
+This is exactly the free operation in imaginarity resource theory, thus it is appropriate to define the free channel by Choi state.
+Let $\mathcal{SC}_{ABA^{\prime}B^{\prime}}$ be the sets of all superchannels $\Theta$ from $\mathcal{C}_{AB}$ to $\mathcal{C}_{A^{\prime}B^{\prime}}$.
+The superchannel $\Theta$ can be represented by Choi matrix [14],
+$J_{\Theta}$ $=$ $\sum_{j,k,\alpha,\beta}|j\alpha\rangle\langle k\beta|\otimes\Theta(|j\alpha\rangle\langle k\beta|)$
+$=$ $\sum_{j,k,\alpha,\beta}\Theta_{j,k,\alpha,\beta,j^{\prime},k^{\prime},\alpha^{\prime},\beta^{\prime}}|j\alpha j^{\prime}\alpha\rangle\langle k\beta k^{\prime}\beta^{\prime}|,$
+where $\Theta_{j,k,\alpha,\beta,j^{\prime},k^{\prime},\alpha^{\prime},\beta^{\prime}}=\langle j^{\prime}\alpha^{\prime}|\Theta(|j\alpha\rangle\langle k\beta|)|k^{\prime}\beta^{\prime}\rangle$ with $|j\alpha\rangle$ means the tensor of the basis from systems $A$ and $B$.
+Similar to the case of channel, superchannel $\Theta$ also has the expression of Kraus operators $\left\{M_{m}\right\}_{m}$ with $\chi_{\Theta(\phi)}=\sum_{m}M_{m}\chi_{\phi}M_{m}^{\dagger}\triangleq\tilde{\Theta}(\chi_{\phi})$, where
+$M_{m}=\sum_{j,j^{\prime},\alpha,\alpha^{\prime}}M_{m,j^{\prime},j,\alpha^{\prime},\alpha}|j^{\prime}\alpha\rangle\langle j\alpha|$
+with $\Theta_{j,k,\alpha,\beta,j^{\prime},k^{\prime},\alpha^{\prime},\beta^{\prime}}=\sum_{m}M_{m,j^{\prime},j,\alpha^{\prime},\alpha}M_{m,k^{\prime},k,\beta^{\prime},\beta}$.
+Here, $\tilde{\Theta}$ means the channel corresponding to the superchannel [14, 26]. We will not distinguish the signs, $\Theta$ or $\tilde{\Theta}$, since it can be judged by the context.
 We call a superchannel $\Theta:\mathcal{C}_{AB}\longrightarrow\mathcal{C}_{A^{\prime}B^{\prime}}$ a real superchannels if $\Theta(\phi)\in\mathcal{RC}_{A^{\prime}B^{\prime}}$ for any channels $\phi\in\mathcal{RC}_{AB}$.
@@
 The free superchannel is defined as the real superchannel.
-In [24], the author presented a framework for quantifying the imaginarity of quantum channels, they introduced the imaginarity measure $C(\phi)\triangleq C(\chi_{\phi})$ of quantum channels should satisfy the following conditions:
+In [24], the author presented a framework for quantifying the imaginarity of quantum channels.
+They introduced the imaginarity measure $C(\phi)\triangleq C(\chi_{\phi})$ of quantum channels should satisfy the following conditions:
 $\bm{(C_{1})}$. Faithfulness: $C(\phi)\geqslant 0$, and $C(\phi)=0$ if and only if $\phi\in\mathcal{RC}_{AB}$;
 $\bm{(C_{2})}$. Monotonicity: $C(\phi)\geqslant C(\Theta(\phi))$ for any real superchannel $\Theta\in\mathcal{RSC}_{ABA^{\prime}B^{\prime}}$.
-Strong monotonicity and convexity often have a great use and a proper quantum resource theory should include them.
-So we add these to the conditions above, and then the imaginarity measure $C$ of quantum channels becomes the more familiar and complete.
+Actually, strong monotonicity and convexity often have a great use and a proper quantum resource theory should include them.
+So we add these to the conditions above, and then the imaginarity measure $C$ of quantum channels becomes more familiar and complete.
 $\bm{(C_{3})}$.
@@
 Property 1.
-Monotonicity under composition with any real channels, that is,
-$(a)$. Left composition: for any $\psi\in\mathcal{RC}_{BD}$ and any $\phi\in\mathcal{C}_{AB}$, we have: $C(\psi\circ\phi)\leqslant C(\phi)$.
-$(b)$. Right composition: for any $\psi\in\mathcal{RC}_{AB}$ and any $\phi\in\mathcal{C}_{BD}$, we have: $C(\phi\circ\psi)\leqslant C(\phi)$.
+Monotonicity under composition with any real channels:
+$(a)$. Left composition: $C(\psi\circ\phi)\leqslant C(\phi)$ for any $\psi\in\mathcal{RC}_{BD}$ and any $\phi\in\mathcal{C}_{AB}$.
+$(b)$. Right composition: $C(\phi\circ\psi)\leqslant C(\phi)$ for any $\psi\in\mathcal{RC}_{AB}$ and any $\phi\in\mathcal{C}_{BD}$.
+$(C)$. Right composition: $C(\phi\circ\psi)=C(\phi)$ for any invertible $\psi\in\mathcal{RC}_{AB}$ and any $\phi\in\mathcal{C}_{BD}$.
 Proof.
-We only prove $(a)$, because $(b)$ is analogous.
+We only prove $(a)$, because $(b)$ is analogous. $(c)$ can be obtained from $(a)$ and $(b)$.
 For $\psi\in\mathcal{RC}_{AB}$ and $\phi\in\mathcal{C}_{BD}$, we consider a superchannel $\Theta_{\psi}$ with $\Theta_{\psi}=\psi\circ\phi$.
-Due to the condition $\bm{(C_{2})}$, we try to explain $\Theta\in\mathcal{RSC}_{ABAD}$.
-In fact, when superchannel $\Theta_{\psi}$ acts on any real channels, it is impossible to generate imaginarity resource since the channel $\psi$ is real.
-We can know that the superchannel $\Theta_{\psi}$ is a real superchannel from the definition.
+When superchannel $\Theta_{\psi}$ acts on any real channels, it is impossible to generate the imaginarity resource since the channel $\psi$ is real.
+It can be seen that the superchannel $\Theta_{\psi}$ is a real superchannel from the definition.
+The result holds due to the condition $\bm{(C_{2})}$.
 ∎
-We can also see that if $\psi$ is invertible, $C(\psi\circ\phi)=C(\phi)$ holds.
 Property 2.
-The combination of imaginarity measures is still an imaginarity measure, that is, for any $\left\{C_{j}\right\}$ and any $\left\{q_{j}\right\}$ with $q_{j}\geqslant 0$, $C\triangleq\sum_{j}q_{j}C_{j}$ is an imaginarity measure .
+The combination of imaginarity measures is still an imaginarity measure, that is, for any $\left\{C_{j}\right\}$ and any probability distribution $\left\{q_{j}\right\}$ with $q_{j}\geqslant 0$, $C\triangleq\sum_{j}q_{j}C_{j}$ is an imaginarity measure.
 Proof.
-It is obvious that $C(\phi)\geqslant 0$ and due to $q_{j}\geqslant 0$, $C(\phi)=0$ is equivalent to every $C_{j}(\phi)=0$, certainly it indicate $\phi\in\mathcal{RC}$.
+It is obvious that $C(\phi)\geqslant 0$.
+Due to $q_{j}\geqslant 0$, $C(\phi)=0$ is equivalent to every $C_{j}(\phi)=0$, and it indicate $\phi\in\mathcal{RC}$.
 For any real superchannel $\Theta$, we have $C(\phi)=\sum_{j}q_{j}C_{j}(\phi)\geqslant\sum_{j}q_{j}C_{j}(\Theta(\phi))=C(\Theta(\phi))$.
 For $\Theta\in\mathcal{RSC}$ with $\Theta(\cdot)=\sum_{m}M_{m}\cdot M_{m}^{\dagger}$, the strong monotonicity holds:
-$C(\phi)=\sum_{j}q_{j}C_{j}(\phi)\geqslant\sum_{j}q_{j}\sum_{m}p_{m}C_{j}(\phi_{m})=\sum_{m}p_{m}\sum_{j}q_{j}C_{j}(\phi_{m})=\sum_{m}p_{m}C(\phi_{m}).$
+$C(\phi)$
+$=$ $\sum_{j}q_{j}C_{j}(\phi)$
+$\geqslant$ $\sum_{j}q_{j}\sum_{m}p_{m}C_{j}(\phi_{m})$
+$=$ $\sum_{m}p_{m}\sum_{j}q_{j}C_{j}(\phi_{m})$
+$=$ $\sum_{m}p_{m}C(\phi_{m}).$
 For any $\left\{\phi_{m}\right\}\subset\mathcal{C}_{AB}$ and any probability distribution $\left\{p_{m}\right\}$, the convexity holds:
-$C(\sum_{m}p_{m}\phi_{m})=\sum_{j}q_{j}C_{j}(\sum_{m}p_{m}\phi_{m})\leqslant\sum_{j}q_{j}\sum_{m}p_{m}C_{j}(\phi_{m})=\sum_{m}p_{m}\sum_{j}q_{j}C_{j}(\phi_{m})=\sum_{m}p_{m}C(\phi_{m}).$
+$C(\sum_{m}p_{m}\phi_{m})$
+$=$ $\sum_{j}q_{j}C_{j}(\sum_{m}p_{m}\phi_{m})$
+$\leqslant$ $\sum_{j}q_{j}\sum_{m}p_{m}C_{j}(\phi_{m})$
+$=$ $\sum_{m}p_{m}\sum_{j}q_{j}C_{j}(\phi_{m})$
+$=$ $\sum_{m}p_{m}C(\phi_{m}).$
 Therefore, $C$ is an imaginarity measure. ∎
 Property 3.
-If $C$ is an imaginarity measure for channel $\phi$, then, quantifier $\hat{C}$ as:
+If $C$ is an imaginarity measure for channel $\phi$, then the quantifier $\hat{C}$ as
 $\hat{C}(\phi)\triangleq\min_{\left\{q_{j},\phi_{j}\right\}}\sum_{j}q_{j}C(\phi_{j}),$
-is a proper measure for $\phi$, where $\min_{\left\{q_{j},\phi_{j}\right\}}$ represents the min of all possible channels decompositions as $\phi=\sum_{j}q_{j}\phi_{j}$ with $q_{j}\geqslant 0$.
+is a proper measure for $\phi$, where $\min_{\left\{q_{j},\phi_{j}\right\}}$ represents the min of all possible channels decompositions as $\phi=\sum_{j}q_{j}\phi_{j}$ with probability distribution $q_{j}\geqslant 0$.
 Proof.
@@
 We can get $C(\phi_{k})=0$ since $q\geqslant 0$, that is, $\phi_{k}\in\mathcal{RC}$, thus $\phi\in\mathcal{RC}$, the faithfulness holds.
-We can notice that, for any $\left\{q_{j}\right\}$, let every $\phi_{j}=\phi$, we can get $\hat{C}(\phi)=\min_{\left\{q_{j},\phi_{j}\right\}}\sum_{j}q_{j}C(\phi_{j})\leqslant C(\phi)$.
+For any probability distribution $\left\{q_{j}\right\}$, if we suppose every $\phi_{j}=\phi$, we can have $\sum_{j}q_{j}C(\phi_{j})=C(\phi)$.
+Thus, it follows that $\hat{C}(\phi)=\min_{\left\{q_{j},\phi_{j}\right\}}\sum_{j}q_{j}C(\phi_{j})\leqslant C(\phi)$.
 For any quantum channel $\phi$, we take the best decomposition of $\phi$, expressed as $\phi=\sum_{k}q_{k}\phi_{k}$, so that $\hat{C}(\phi)=\sum_{k}q_{k}C(\phi_{k})$.
 For $\Theta\in\mathcal{RSC}$ with $\Theta(\cdot)=\sum_{m}M_{m}\cdot M_{m}^{\dagger}$, we denote $a_{k,m}=\text{Tr}(M_{m}\chi_{\phi_{k}}M_{m}^{\dagger})$, $p_{m}=\text{Tr}(M_{m}\chi_{\phi}M_{m}^{\dagger})$, then
-$\hat{C}(\phi)=\sum_{k}q_{k}C(\phi_{k})$$\geqslant$ $\sum_{k}q_{k}\sum_{m}a_{k,m}C(\frac{M_{m}\chi_{\phi_{k}}M_{m}^{\dagger}}{a_{k,m}})=\sum_{m}p_{m}\sum_{k}\frac{q_{k}a_{k,m}}{p_{m}}C(\frac{M_{m}\chi_{\phi_{k}}M_{m}^{\dagger}}{a_{k,m}})$
-$\geqslant$ $\sum_{m}p_{m}\sum_{k}C(\frac{q_{k}a_{k,m}}{p_{m}}\frac{M_{m}\chi_{\phi_{k}}M_{m}^{\dagger}}{a_{k,m}})=\sum_{m}p_{m}C(\frac{\sum_{k}q_{k}M_{m}\chi_{\phi_{k}}M_{m}^{\dagger}}{p_{m}})$
-$=$ $\sum_{m}p_{m}C(\frac{M_{m}\chi_{\phi}M_{m}^{\dagger}}{p_{m}})\geqslant\sum_{m}p_{m}\hat{C}(\frac{M_{m}\chi_{\phi}M_{m}^{\dagger}}{p_{m}}).$
-The first inequality is due to the strong monotonicity of $C$, the second inequality is due to the convexity of $C$ and $\sum_{k}\frac{q_{k}a_{k,m}}{p_{m}}=\frac{\sum_{k}q_{k}\text{Tr}(M_{m}\chi_{\phi_{k}}M_{m}^{\dagger})}{\text{Tr}(M_{m}\chi_{\phi}M_{m}^{\dagger})}=1$.
-To verify convexity of $\hat{C}$ for simplicity, we only consider two quantum channels $\phi_{1}$ and $\phi_{2}$, and define $\psi=a\phi_{1}+b\phi_{2}$ with $a+b=1$.
-We take the best decomposition of $\phi_{1}$ and $\phi_{2}$ respectively as $\phi_{1}=\sum_{m}q_{1,m}\phi_{1,m}$ and $\phi_{2}=\sum_{n}q_{2,n}\phi_{2,n}$ so that the equations hold:
+$\hat{C}(\phi)$
+$=$ $\sum_{k}q_{k}C(\phi_{k})$
+$\geqslant$ $\sum_{k}q_{k}\sum_{m}a_{k,m}C(\frac{M_{m}\chi_{\phi_{k}}M_{m}^{\dagger}}{a_{k,m}})$
+$=$ $\sum_{m}p_{m}\sum_{k}\frac{q_{k}a_{k,m}}{p_{m}}C(\frac{M_{m}\chi_{\phi_{k}}M_{m}^{\dagger}}{a_{k,m}})$
+$\geqslant$ $\sum_{m}p_{m}C(\sum_{k}\frac{q_{k}a_{k,m}}{p_{m}}\frac{M_{m}\chi_{\phi_{k}}M_{m}^{\dagger}}{a_{k,m}})$
+$=$ $\sum_{m}p_{m}C(\frac{\sum_{k}q_{k}M_{m}\chi_{\phi_{k}}M_{m}^{\dagger}}{p_{m}})$
+$=$ $\sum_{m}p_{m}C(\frac{M_{m}\chi_{\phi}M_{m}^{\dagger}}{p_{m}})$
+$\geqslant$ $\sum_{m}p_{m}\hat{C}(\frac{M_{m}\chi_{\phi}M_{m}^{\dagger}}{p_{m}}).$
+The first inequality follows from the strong monotonicity of $C$.
+The second inequality arises from the convexity of $C$ and the fact that $\sum_{k}\frac{q_{k}a_{k,m}}{p_{m}}=\frac{\sum_{k}q_{k}\text{Tr}(M_{m}\chi_{\phi_{k}}M_{m}^{\dagger})}{\text{Tr}(M_{m}\chi_{\phi}M_{m}^{\dagger})}=1$.
+The third inequality results from $\hat{C}(\phi)\leqslant C(\phi)$.
+To verify the convexity of $\hat{C}$, for simplicity and without loss of generality, we consider only two quantum channels $\phi_{1}$ and $\phi_{2}$.
+We define $\psi=a\phi_{1}+b\phi_{2}$ with $a+b=1$.
+We take the best decomposition of $\phi_{1}$ and $\phi_{2}$ respectively as $\phi_{1}=\sum_{m}q_{1,m}\phi_{1,m}$ and $\phi_{2}=\sum_{n}q_{2,n}\phi_{2,n}$ so that the following equations hold:
 $\begin{cases}\hat{C}(\phi_{1})=\sum_{m}q_{1,m}C(\phi_{1,m}),\\ \hat{C}(\phi_{2})=\sum_{n}q_{2,n}C(\phi_{2,n}).\end{cases}$
 Then, $a\hat{C}(\phi_{1})+b\hat{C}(\phi_{2})=\sum_{m}aq_{1,m}C(\phi_{1,m})+\sum_{n}bq_{2,n}C(\phi_{2,n})$.
-For ease of writing, if $m\neq n$, we supplement $\left\{0,\bm{0}\right\}$ for the less set so that both sets have the same quantity, that is, $a\hat{C}(\phi_{1})+b\hat{C}(\phi_{2})=\sum_{m}aq_{1,m}C(\phi_{1,m})+bq_{2,m}C(\phi_{2,m})$.
-Substitute the new symbol $\left\{q_{2m-1},\psi_{2m-1}\right\}$ and $\left\{q_{2m},\psi_{2m}\right\}$ for $\left\{aq_{1,m},\phi_{1,m}\right\}$ and $\left\{bq_{2,m},\phi_{2,m}\right\}$ respectively.
-We have
-$a\hat{C}(\phi_{1})+b\hat{C}(\phi_{2})=\sum_{m}q_{2m-1}C(\psi_{2m-1})+q_{2m}C(\psi_{2m})\geqslant\min_{q_{j},\psi_{j}}\sum_{j}q_{j}C(\psi_{j})=\hat{C}(\psi)=\hat{C}(a\phi_{1}+b\phi_{2}).$
-In this way, it can be known that $\hat{C}$ is a proper imaginarity measure. ∎
+For ease of writing, if $m\neq n$, we supplement the elements $\left\{0,\bm{0}\right\}$ to the smaller set so that both sets have the same number of elements.
+This gives $a\hat{C}(\phi_{1})+b\hat{C}(\phi_{2})=\sum_{m}aq_{1,m}C(\phi_{1,m})+bq_{2,m}C(\phi_{2,m})$.
+Next, we introduce the new symbols $\left\{q_{2m-1},\psi_{2m-1}\right\}$ and $\left\{q_{2m},\psi_{2m}\right\}$ to represent $\left\{aq_{1,m},\phi_{1,m}\right\}$ and $\left\{bq_{2,m},\phi_{2,m}\right\}$ respectively.
+Then we have
+$a\hat{C}(\phi_{1})+b\hat{C}(\phi_{2})$
+$=$ $\sum_{m}q_{2m-1}C(\psi_{2m-1})+q_{2m}C(\psi_{2m})$
+$\geqslant$ $\min_{q_{j},\psi_{j}}\sum_{j}q_{j}C(\psi_{j})$
+$=$ $\hat{C}(\psi)=\hat{C}(a\phi_{1}+b\phi_{2}).$
+In this way, we confirm that $\hat{C}$ is a valid imaginarity measure. ∎
 3 Alternative framework for the imaginarity of quantum channels
-In the previous section we refined the framework.
-However, the condition $\bm{(C_{3})}$ and $\bm{(C_{4})}$ are difficult to verify in some cases.
-Here, inspired by the similar alternative framework in many resource theory [22, 23], we have the following result.
+In Section 2, we introduce the framework.
+However, conditions $\bm{(C_{3})}$ and $\bm{(C_{4})}$ can sometimes be difficult to verify.
+In coherence resource theory, to simplify the process of verifying whether a quantifier qualifies as a coherence measure, an alternative framework for quantifying coherence has been proposed [22].
+Similar alternative frameworks can also be established in many other resource theories, such as imaginarity resource theory [7], and coherence of quantum channels [23].
+Inspired by [23], we Introduce the following condition:
+$\bm{(C_{5})}$.
+Additivity: $C(\phi)=p_{1}C(\phi_{1})+p_{2}C(\phi_{2})$ for any $p_{1}$, $p_{2}$ satisfied $p_{1}+p_{2}=1$, any $\phi_{1}\in\mathcal{C}_{AB_{1}}$ and $\phi_{2}\in\mathcal{C}_{AB_{2}}$, where $\phi=p_{1}\phi_{1}\oplus p_{2}\phi_{2}$, $\phi\in\mathcal{C}_{AB}$ with $|B|=|B_{1}|+|B_{2}|$.
+When conditions $\bm{(C_{1})}$ and $\bm{(C_{2})}$ hold, the conditions $\bm{(C_{3})}$ and $\bm{(C_{4})}$ are equivalent to the condition $\bm{(C_{5})}$.
+To prove this, we first state the following theorem:
 Theorem 1.
-$\bm{(C_{3})}+\bm{(C_{4})}$ is equivalent to $\bm{(C_{5})}$ when condition $\bm{(C_{1})}$ and $\bm{(C_{2})}$ holds. The condition $\bm{(C_{5})}$ is:
-$\bm{(C_{5})}$.
-Additivity: For any $p_{1}$, $p_{2}$ satisfied $p_{1}+p_{2}=1$, any $\phi_{1}\in\mathcal{C}_{AB_{1}}$ and $\phi_{2}\in\mathcal{C}_{AB_{2}}$, let $\phi=p_{1}\phi_{1}\oplus p_{2}\phi_{2}$, $\phi\in\mathcal{C}_{AB}$ with $|B|=|B_{1}|+|B_{2}|$, then $C(\phi)=p_{1}C(\phi_{1})+p_{2}C(\phi_{2})$ holds.
+$\bm{(C_{3})}$ and $\bm{(C_{4})}$ give rise to $\bm{(C_{5})}$ when condition $\bm{(C_{1})}$ and $\bm{(C_{2})}$ holds.
 Proof.
-The proof is long and we divide it into steps.
-$\bm{(C_{3})}+\bm{(C_{4})}\Rightarrow\bm{(C_{5})}$:
-(STEP 1).
+STEP 1, obtain: $\bm{C(\phi)\geqslant p_{1}C(\tilde{\phi}_{1})+p_{2}C(\tilde{\phi}_{2})}$.
 Let $Q_{1}=|0\rangle\langle 0|+\cdots+||B_{1}|-1\rangle\langle|B_{1}|-1|+||B|\rangle\langle|B||+\cdots+||B|+|B_{1}|-1\rangle\langle|B|+|B_{1}|-1|+|(|A|-1)|B|\rangle\langle(|A|-1)|B||+\cdots+|(|A|-1)|B|+|B_{1}|-1\rangle\langle(|A|-1)|B|+|B_{1}|-1|$,
@@
 We can give an intuitive expression.
-$Q_{1}=\overbrace{\begin{pmatrix}I_{|B_{1}|}&&&&\\ &O_{|B_{2}|}&&&\\ &&\ddots&&\\ &&&I_{|B_{1}|}&\\ &&&&O_{|B_{2}|}\end{pmatrix}}^{|A||B|}=\bigoplus\limits_{j=0}^{|A|-1}\begin{pmatrix}I_{|B_{1}|}&\\ &O_{|B_{2}|}\end{pmatrix},$
-$Q_{2}=\underbrace{\begin{pmatrix}O_{|B_{1}|}&&&&\\ &I_{|B_{2}|}&&&\\ &&\ddots&&\\ &&&O_{|B_{1}|}&\\ &&&&I_{|B_{2}|}\end{pmatrix}}_{|A||B|}=\bigoplus\limits_{j=0}^{|A|-1}\begin{pmatrix}O_{|B_{1}|}&\\ &I_{|B_{2}|}\end{pmatrix},$
+$Q_{1}=\overbrace{\begin{pmatrix}I_{|B_{1}|}&&&&\\ &O_{|B_{2}|}&&&\\ &&\ddots&&\\ &&&I_{|B_{1}|}&\\ &&&&O_{|B_{2}|}\end{pmatrix}}^{|A||B|}$
+$=\bigoplus\limits_{j=0}^{|A|-1}\begin{pmatrix}I_{|B_{1}|}&\\ &O_{|B_{2}|}\end{pmatrix},$
+$Q_{2}=\underbrace{\begin{pmatrix}O_{|B_{1}|}&&&&\\ &I_{|B_{2}|}&&&\\ &&\ddots&&\\ &&&O_{|B_{1}|}&\\ &&&&I_{|B_{2}|}\end{pmatrix}}_{|A||B|}$
+$=\bigoplus\limits_{j=0}^{|A|-1}\begin{pmatrix}O_{|B_{1}|}&\\ &I_{|B_{2}|}\end{pmatrix},$
 where $I_{a}$ and $O_{b}$ denote identity matrix and zero matrix with dimensions $a$ and $b$ respectively.
-Due to $Q_{1}^{\dagger}Q_{1}+Q_{2}^{\dagger}Q_{2}=I_{|A||B|}$ and $Q_{1}Q_{2}=O_{|A||B|}$, let $\Theta(\cdot)=Q_{1}\cdot Q_{1}^{\dagger}+Q_{2}\cdot Q_{2}^{\dagger}$, we can know that $\Theta\in\mathcal{RSC}_{ABA^{\prime}B^{\prime}}$.
+Given that $Q_{1}^{\dagger}Q_{1}+Q_{2}^{\dagger}Q_{2}=I_{|A||B|}$ and $Q_{1}Q_{2}=O_{|A||B|}$, let $\Theta(\cdot)=Q_{1}\cdot Q_{1}^{\dagger}+Q_{2}\cdot Q_{2}^{\dagger}$.
+It follows that $\Theta\in\mathcal{RSC}_{ABAB}$.
 Under the requirement of $\bm{(C_{5})}$, the Choi state of $\phi$ is
-$\chi_{\phi}=\sum_{j,k}|j\rangle\langle k|\otimes\phi(|j\rangle\langle k|)/|A|=\sum_{j,k}|j\rangle\langle k|\otimes\big{[}p_{1}\phi_{1}(|j\rangle\langle k|)\oplus p_{2}\phi_{2}(|j\rangle\langle k|\big{]}/|A|.$
+$\chi_{\phi}$
+$=$ $\sum_{j,k}|j\rangle\langle k|\otimes\phi(|j\rangle\langle k|)/|A|$
+$=$ $\sum_{j,k}|j\rangle\langle k|\otimes\big{[}p_{1}\phi_{1}(|j\rangle\langle k|)\oplus p_{2}\phi_{2}(|j\rangle\langle k|\big{]}/|A|.$
 Thus,
 $Q_{1}\chi_{\phi}Q_{1}^{\dagger}=\sum_{j,k}|j\rangle\langle k|\otimes\big{[}p_{1}\phi_{1}(|j\rangle\langle k|)\oplus\bm{0}_{2}(|j\rangle\langle k|\big{]}/|A|.$
-It’s obvious that $\text{Tr}(Q_{1}\chi_{\phi}Q_{1}^{\dagger})$ is exactly $p_{1}$.
+It is clear that $\text{Tr}(Q_{1}\chi_{\phi}Q_{1}^{\dagger})$ is exactly $p_{1}$.
 Let $\chi_{\tilde{\phi}_{1}}=Q_{1}\chi_{\phi}Q_{1}^{\dagger}/p_{1}=\sum_{j,k}|j\rangle\langle k|\otimes\big{[}\phi_{1}(|j\rangle\langle k|)\oplus\bm{0}_{2}(|j\rangle\langle k|\big{]}/|A|$.
-Similarly we obtain $\chi_{\tilde{\phi}_{2}}$.
-At this point, the requirement of $\bm{(C_{3})}$ has been satisfied, that is
-$\begin{cases}\Theta\in\mathcal{RSC}_{ABA^{\prime}B^{\prime}}\text{ with }\Theta(\cdot)=Q_{1}\cdot Q_{1}^{\dagger}+Q_{2}\cdot Q_{2}^{\dagger},\\ p_{m}=\text{Tr}(Q_{m}\chi_{\phi}Q_{m}^{\dagger}),\chi_{\tilde{\phi}_{m}}=Q_{m}\chi_{\phi}Q_{m}^{\dagger}/p_{m},\end{cases}$
-we have
-$C(\chi_{\phi})\geqslant p_{1}C(\chi_{\tilde{\phi}_{1}})+p_{2}C(\chi_{\tilde{\phi}_{2}}).$ (1)
-(STEP 2).
+Similarly we can obtain $\chi_{\tilde{\phi}_{2}}$.
+From the condition $\bm{(C_{3})}$, we have:
+$C(\phi)=C(\chi_{\phi})$$\geqslant$ $p_{1}C(\chi_{\tilde{\phi}_{1}})+p_{2}C(\chi_{\tilde{\phi}_{2}})$ (1)
+$=$ $p_{1}C(\tilde{\phi}_{1})+p_{2}C(\tilde{\phi}_{2}).$
+STEP 2, obtain: $\bm{C(\phi)\leqslant p_{1}C(\tilde{\phi}_{1})+p_{2}C(\tilde{\phi}_{2})}$.
 For any $\rho\in\mathcal{D}_{A}$, $\rho=\sum_{j,k}\rho_{j,k}|j\rangle\langle k|$, any $\phi\in\mathcal{C}_{AB}$, $\phi(\rho)$ can be represented by $J_{\phi}$, in order to do this, $\rho$ and $J_{\phi}$ need to be tensor index realigned:
 $\begin{cases}J^{\prime}_{\phi}=\sum_{j,k}\langle jk|\otimes\phi(|j\rangle\langle k|),\\ \rho^{\prime}=\sum_{j,k}\rho_{j,k}|jk\rangle,\end{cases}$
-then $\phi(\rho)=J^{\prime}_{\phi}\rho^{\prime}=\sum_{j,k}\rho_{j,k}\phi(|j\rangle\langle k|)$. Under the requirement of $(C_{5})$, we have
-$\phi(\rho)$ $=$ $\sum_{j,k}\rho_{j,k}\phi(|j\rangle\langle k|)=\sum_{j,k}\rho_{j,k}\big{[}p_{1}\phi_{1}(|j\rangle\langle k|)\oplus p_{2}\phi_{2}(|j\rangle\langle k|)\big{]}$
-$=$ $p_{1}\sum_{j,k}\rho_{j,k}\big{[}\phi_{1}(|j\rangle\langle k|)\oplus\bm{0}_{2}(|j\rangle\langle k|)\big{]}+p_{2}\sum_{j,k}\rho_{j,k}\big{[}\bm{0}_{1}(|j\rangle\langle k|)\oplus\phi_{2}(|j\rangle\langle k|)\big{]}$
+then $\phi(\rho)=J^{\prime}_{\phi}\rho^{\prime}=\sum_{j,k}\rho_{j,k}\phi(|j\rangle\langle k|)$. Under the requirement of $(C_{5})$, for any $\rho$ we have
+$\phi(\rho)$
+$=$ $\sum_{j,k}\rho_{j,k}\phi(|j\rangle\langle k|)$
+$=$ $\sum_{j,k}\rho_{j,k}\big{[}p_{1}\phi_{1}(|j\rangle\langle k|)\oplus p_{2}\phi_{2}(|j\rangle\langle k|)\big{]}$
+$=$ $p_{1}\sum_{j,k}\rho_{j,k}\big{[}\phi_{1}(|j\rangle\langle k|)\oplus\bm{0}_{2}(|j\rangle\langle k|)\big{]}$
+$+p_{2}\sum_{j,k}\rho_{j,k}\big{[}\bm{0}_{1}(|j\rangle\langle k|)\oplus\phi_{2}(|j\rangle\langle k|)\big{]}$
 $=$ $p_{1}\tilde{\phi}_{1}(\rho)+p_{2}\tilde{\phi}_{2}(\rho),$
-that is $\phi=p_{1}\tilde{\phi}_{1}+p_{2}\tilde{\phi}_{2}$. From $\bm{(C_{4})}$ We can know that
+that is $\phi=p_{1}\tilde{\phi}_{1}+p_{2}\tilde{\phi}_{2}$. From $\bm{(C_{4})}$, it follows that
 $C(\phi)\leqslant p_{1}C(\tilde{\phi}_{1})+p_{2}C(\tilde{\phi}_{2}).$ (2)
-(STEP 3).
-From (1) in (STEP 1) and (2) in (STEP 2), we obtain $C(\phi)=p_{1}C(\tilde{\phi}_{1})+p_{2}C(\tilde{\phi}_{2})$.
-Next we need to prove $C(\phi_{1})=C(\tilde{\phi}_{1})$, that is, we need to prove $C(\chi_{\phi_{1}})=C(\chi_{\tilde{\phi}_{1}})$.
-First, we specify the range of index $i$ and $j$, let $i=0,1,\cdots,|B|-1$, $j=0,1,\cdots,|B_{1}|-1$.
+STEP 3, obtain $\bm{(C_{5})}$ since $\bm{C(\phi_{j})=C(\tilde{\phi}_{j})}$, $\bm{j=1,2}$.
+From inequalities (1) and (2) above, we obtain $C(\phi)=p_{1}C(\tilde{\phi}_{1})+p_{2}C(\tilde{\phi}_{2})$.
+Next, we need to prove that $C(\phi_{1})=C(\tilde{\phi}_{1})$, which requires us to show that $C(\chi_{\phi_{1}})=C(\chi_{\tilde{\phi}_{1}})$.
+First, we specify the range of indices $i$ and $j$, let $i=0,1,\cdots,|B|-1$, $j=0,1,\cdots,|B_{1}|-1$.
 Let $K_{0}^{1}$ satisfies $\langle i|K_{0}^{1}|j\rangle=\delta_{i,j}$.
-An intuitive expression of $K_{0}^{1}$ is
+An intuitive expression for $K_{0}^{1}$ is
 $K_{0}^{1}=\left.\begin{pmatrix}I_{|B_{1}|}\\ \hat{O}\end{pmatrix}\right\}|B|,$
-where $\hat{O}$ denotes zero matrix with dimension $|B_{2}|\times|B_{1}|$.
-Let $K_{n}^{2}$ satisfy $\langle j|K_{n}^{2}|i\rangle=\delta_{i,j+n|B_{1}|}$, $n=0,1,\cdots,\left\lceil|B_{2}|/|B_{1}|\right\rceil$.
-$\left\lceil a\right\rceil$ is the ceil of $a$.
-We can give an example to make this easier to understand.
-See Example 1 after this proof.
+where $\hat{O}$ denotes the zero matrix with dimension $|B_{2}|\times|B_{1}|$.
+Let $K_{n}^{2}$ satisfy $\langle j|K_{n}^{2}|i\rangle=\delta_{i,j+n|B_{1}|}$ for $n=0,1,\cdots,$ $\left\lceil|B_{2}|/|B_{1}|\right\rceil$, where $\left\lceil a\right\rceil$ means the ceil of $a$.
 Let
-$\mathcal{K}_{0}^{1}=\bigoplus_{j=0}^{|A|-1}K_{0}^{1}=\begin{pmatrix}K_{0}^{1}&&\\ &\ddots&\\ &&K_{0}^{1}\end{pmatrix},\mathcal{K}_{n}^{2}=\bigoplus_{j=0}^{|A|-1}K_{n}^{2}=\begin{pmatrix}K_{n}^{2}&&\\ &\ddots&\\ &&K_{n}^{2}\end{pmatrix},$
-we can verify $(\mathcal{K}_{0}^{1})^{\dagger}\mathcal{K}_{0}^{1}=I_{|B_{1}|}$ and $\sum_{j=0}^{\left\lceil|B_{2}|/|B_{1}|\right\rceil}(\mathcal{K}_{n}^{2})^{\dagger}\mathcal{K}_{n}^{2}=I_{|B|}$.
+$\mathcal{K}_{0}^{1}=\bigoplus_{j=0}^{|A|-1}K_{0}^{1}=\begin{pmatrix}K_{0}^{1}&&\\ &\ddots&\\ &&K_{0}^{1}\end{pmatrix},$
+$\mathcal{K}_{n}^{2}=\bigoplus_{j=0}^{|A|-1}K_{n}^{2}=\begin{pmatrix}K_{n}^{2}&&\\ &\ddots&\\ &&K_{n}^{2}\end{pmatrix}.$
+We have $(\mathcal{K}_{0}^{1})^{\dagger}\mathcal{K}_{0}^{1}=I_{|B_{1}|}$ and $\sum_{j=0}^{\left\lceil|B_{2}|/|B_{1}|\right\rceil}(\mathcal{K}_{n}^{2})^{\dagger}\mathcal{K}_{n}^{2}=I_{|B|}$.
 Thus $\Theta_{1}\in\mathcal{RSC}_{AB_{1}AB}$ with $\Theta_{1}(\cdot)=\mathcal{K}_{0}^{1}\cdot(\mathcal{K}_{0}^{1})^{\dagger}$, $\Theta_{2}\in\mathcal{RSC}_{ABAB_{1}}$ with $\Theta_{2}(\cdot)=\sum_{j=0}^{\left\lceil|B_{2}|/|B_{1}|\right\rceil}\mathcal{K}_{n}^{2}\cdot(\mathcal{K}_{n}^{2})^{\dagger}$.
-We can also obtain $\Theta_{1}(\chi_{\phi_{1}})=\chi_{\tilde{\phi}_{1}}$ and $\Theta_{2}(\chi_{\tilde{\phi}_{1}})=\chi_{\phi_{1}}$.
+It can be seen that $\Theta_{1}(\chi_{\phi_{1}})=\chi_{\tilde{\phi}_{1}}$ and $\Theta_{2}(\chi_{\tilde{\phi}_{1}})=\chi_{\phi_{1}}$.
 Combining with $\bm{(C_{2})}$, the above implies that the following inequality holds:
-$C(\chi_{\phi_{1}})=C(\Theta_{2}(\chi_{\tilde{\phi}_{1}}))\leqslant C(\chi_{\tilde{\phi}_{1}})=C(\Theta_{1}(\chi_{\phi_{1}}))\leqslant C(\chi_{\phi_{1}}).$
+$C(\chi_{\phi_{1}})$
+$=$ $C(\Theta_{2}(\chi_{\tilde{\phi}_{1}}))$
+$\leqslant$ $C(\chi_{\tilde{\phi}_{1}})$
+$=$ $C(\Theta_{1}(\chi_{\phi_{1}}))$
+$\leqslant$ $C(\chi_{\phi_{1}}).$
 As a result, $C(\chi_{\phi_{1}})=C(\chi_{\tilde{\phi}_{1}})$. It follows from a similar discussion that $C(\chi_{\phi_{2}})=C(\chi_{\tilde{\phi}_{2}})$.
-Up to now, we prove $\bm{(C_{5})}$ holds: $C(\phi)=p_{1}C(\phi_{1})+p_{2}C(\phi_{2})$.
-$\bm{(C_{5})}\Rightarrow\bm{(C_{3})}+\bm{(C_{4})}$:
-(STEP 4).
+Up to now, we prove $\bm{(C_{5})}$ holds: $C(\phi)=p_{1}C(\phi_{1})+p_{2}C(\phi_{2})$. ∎
+Under the requirement of $\bm{(C_{5})}$, we have $\chi_{\phi_{1}}$ and $\chi_{\phi_{2}}$ below.
+$\begin{cases}\chi_{\phi_{1}}=\sum_{j,k}|j\rangle\langle k|\otimes\phi_{1}(|j\rangle\langle k|)/|A|,\\ \chi_{\phi_{2}}=\sum_{j,k}|j\rangle\langle k|\otimes\phi_{2}(|j\rangle\langle k|)/|A|.\end{cases}$
+We also define $\chi_{\tilde{\phi}_{1}}$ and $\chi_{\tilde{\phi}_{2}}$ below.
+$\begin{cases}\chi_{\tilde{\phi}_{1}}=\sum_{j,k}|j\rangle\langle k|\otimes\big{[}\phi_{1}(|j\rangle\langle k|)\oplus\bm{0}_{2}(|j\rangle\langle k|\big{]}/|A|,\\ \chi_{\tilde{\phi}_{2}}=\sum_{j,k}|j\rangle\langle k|\otimes\big{[}\bm{0}_{1}(|j\rangle\langle k|)\oplus\phi_{2}(|j\rangle\langle k|\big{]}/|A|.\end{cases}$
+Thus the following property holds.
+Next, we prove the reverse result:
+Theorem 2.
+$\bm{(C_{5})}$ gives rise to $\bm{(C_{3})}$ and $\bm{(C_{4})}$ when condition $\bm{(C_{1})}$ and $\bm{(C_{2})}$ holds.
+Proof.
+prove $\bm{(C_{3})}$.
 For $\Theta\in\mathcal{RSC}_{ABA^{\prime}B^{\prime}}$ with $\Theta(\cdot)=\sum_{m}M_{m}\cdot M_{m}^{\dagger}$, we introduce an auxiliary system $S$ with dimension $N$ and orthonormal basis $\left\{|n\rangle\right\}_{n=0}^{N-1}$.
-The system $AB$ formed by Choi state and the auxiliary system $S$ form a combined system $ABS$, of which the basis is $\left\{|j\alpha\rangle\otimes|n\rangle\right\}$ and the whole system is initially in the state $\chi_{\phi}\otimes|0\rangle\langle 0|\triangleq\chi_{\phi}^{ABS}$.
+The system $AB$, formed by Choi state and the auxiliary system $S$, constitutes a combined system $ABS$, whose basis is $\left\{|j\alpha\rangle\otimes|n\rangle\right\}$.
+The entire system is initially in the state $\chi_{\phi}\otimes|0\rangle\langle 0|\triangleq\chi_{\phi}^{ABS}$.
 Let $U_{n}=\sum_{k=0}^{N-1}|(k+n)\text{ mod }N\rangle\langle k|$, then $\Theta^{ABS}(\cdot)\in\mathcal{RSC}_{ABSA^{\prime}B^{\prime}S}$ with $\Theta^{ABS}(\cdot)=\sum_{n}(M_{n}\otimes U_{n})\cdot(M_{n}\otimes U_{n})^{\dagger}$.
 Thus
-$\Theta^{ABS}(\chi_{\phi}^{ABS})=\sum_{n}(M_{n}\otimes U_{n})(\chi_{\phi}\otimes|0\rangle\langle 0|)(M_{n}\otimes U_{n})^{\dagger}=\sum_{n}M_{n}\chi_{\phi}M_{n}^{\dagger}\otimes|n\rangle\langle n|\triangleq\sum_{n}p_{n}\chi_{\phi_{n}}\otimes|n\rangle\langle n|,$
+$\Theta^{ABS}(\chi_{\phi}^{ABS})$
+$=$ $\sum_{n}(M_{n}\otimes U_{n})(\chi_{\phi}\otimes|0\rangle\langle 0|)(M_{n}\otimes U_{n})^{\dagger}$
+$=$ $\sum_{n}M_{n}\chi_{\phi}M_{n}^{\dagger}\otimes|n\rangle\langle n|$
+$=$ $\sum_{n}p_{n}\chi_{\phi_{n}}\otimes|n\rangle\langle n|,$
 where $p_{n}=\text{Tr}(M_{n}\chi_{\phi}M_{n}^{\dagger})$ and $\chi_{\phi_{n}}=M_{n}\chi_{\phi}M_{n}^{\dagger}/p_{n}$.
-We can know that $\sum_{n}p_{n}=1$, $\phi_{n}\in\mathcal{C}_{AB_{n}}$ with $|B|=\sum_{n}|B_{n}|$. From $\bm{(C_{5})}$, we get
-$\begin{cases}C(\chi_{\phi}^{ABS})=C(\chi_{\phi}\otimes|0\rangle\langle 0|)=C(\chi_{\phi}),\\ C(\Theta^{ABS}(\chi_{\phi}^{ABS}))=C(\sum_{n}p_{n}\chi_{\phi_{n}}\otimes|n\rangle\langle n|)=C(\bigoplus_{n}p_{n}\chi_{\phi_{n}})=\sum_{n}p_{n}C(\chi_{\phi_{n}}).\end{cases}$
-From $\bm{(C_{2})}$, we can obtain
-$C(\chi_{\phi})=C(\chi_{\phi}^{ABS})\geqslant C(\Theta^{ABS}(\chi_{\phi}^{ABS}))=\sum_{n}p_{n}C(\chi_{\phi_{n}}),$
+It can be seen that $\sum_{n}p_{n}=1$, $\phi_{n}\in\mathcal{C}_{AB_{n}}$ with $|B|=\sum_{n}|B_{n}|$. From $\bm{(C_{5})}$, we have
+$C(\chi_{\phi}^{ABS})=C(\chi_{\phi}\otimes|0\rangle\langle 0|)=C(\chi_{\phi}),$
+and
+$C(\Theta^{ABS}(\chi_{\phi}^{ABS}))$
+$=$ $C(\sum_{n}p_{n}\chi_{\phi_{n}}\otimes|n\rangle\langle n|)$
+$=$ $C(\bigoplus_{n}p_{n}\chi_{\phi_{n}})$
+$=$ $\sum_{n}p_{n}C(\chi_{\phi_{n}}).$
+From $\bm{(C_{2})}$ and above, we can obtain
+$C(\chi_{\phi})\geqslant\sum_{n}p_{n}C(\chi_{\phi_{n}}),$
 that is, condition $\bm{(C_{3})}$ holds.
-(STEP 5).
-For any $\left\{\phi_{m}\right\}\subset\mathcal{C}_{AB}$ and any probability distribution $\left\{p_{m}\right\}$, we construct the similar auxiliary system $S$ with orthonormal basis $\left\{|m\rangle\right\}_{m}$ and define $\chi_{\phi}^{ABS}=\sum_{m}p_{m}\chi_{\phi_{m}}\otimes|m\rangle\langle m|$.
-Due to $\sum_{m}(I_{|A||B|}\otimes|0\rangle\langle m|)^{\dagger}(I_{|A||B|}\otimes|0\rangle\langle m|)=I_{|A||B||S|}$, $\Theta^{ABS}\in\mathcal{RSC}_{ABSA^{\prime}B^{\prime}S}$ with $\Theta^{ABS}(\cdot)=\sum_{m}(I_{|A||B|}\otimes|0\rangle\langle m|)\cdot(I_{|A||B|}\otimes|0\rangle\langle m|)^{\dagger}$.
-Then
-$\Theta^{ABS}(\chi_{\phi}^{ABS})=\sum_{m}(I_{|A||B|}\otimes|0\rangle\langle m|)\chi_{\phi}^{ABS}(I_{|A||B|}\otimes|0\rangle\langle m|)^{\dagger}=\sum_{m}p_{m}\chi_{\phi_{m}}\otimes|0\rangle\langle 0|.$
-We know from $\bm{(C_{2})}$ and $\bm{(C_{5})}$ that the following inequality holds:
-$\sum_{m}p_{m}C(\chi_{\phi_{m}})=C(\chi_{\phi}^{ABS})\geqslant C(\Theta^{ABS}(\chi_{\phi}^{ABS}))=C(\sum_{m}p_{m}\chi_{\phi_{m}}),$
-that is, condition $\bm{(C_{4})}$ holds.
-Therefore, the proof is finished. ∎
-Example 1.
-Suppose $|B_{1}|=2$ and $|B_{2}|=3$, then $i=0,1,\cdots,4$, $j=0,1$, $n=0,1,2$ and
-$K_{0}^{2}=\begin{pmatrix}1&0&0&0&0\\ 0&1&0&0&0\end{pmatrix}\triangleq\begin{pmatrix}I_{|B_{1}|}&O_{|B_{1}|}&\tilde{O}\end{pmatrix},$
-$K_{1}^{2}=\begin{pmatrix}0&0&1&0&0\\ 0&0&0&1&0\end{pmatrix}\triangleq\begin{pmatrix}O_{|B_{1}|}&I_{|B_{1}|}&\tilde{O}\end{pmatrix},$
-$K_{2}^{2}=\begin{pmatrix}0&0&0&0&1\\ 0&0&0&0&0\end{pmatrix}\triangleq\begin{pmatrix}O_{|B_{1}|}&O_{|B_{1}|}&\tilde{I}\end{pmatrix},$
-where $\tilde{O}$ denotes the first $(|B_{2}|\text{ mod }|B_{1}|)$ columns of $O_{|B_{1}|}$, same does $\tilde{I}$.
-It is often difficult to verify whether a measure has strong monotonicity and convexity directly, which may be easier after using Theorem 1.
-When verifying whether a quantifier is a proper measure, we can choose verifying $\bm{(C_{1})}$$\bm{(C_{4})}$ or verifying $\bm{(C_{1})}+\bm{(C_{2})}+\bm{(C_{5})}$.
-Under the requirement of $\bm{(C_{5})}$, that is, for any $p_{1}$, $p_{2}$ satisfied $p_{1}+p_{2}=1$, any $\phi_{1}\in\mathcal{C}_{AB_{1}}$ and $\phi_{2}\in\mathcal{C}_{AB_{2}}$, let $\phi=p_{1}\phi_{1}\oplus p_{2}\phi_{2}$, $\phi\in\mathcal{C}_{AB}$ with $|B|=|B_{1}|+|B_{2}|$.
-We have $\chi_{\phi_{1}}$ and $\chi_{\phi_{2}}$, and we define $\chi_{\tilde{\phi}_{1}}$ and $\chi_{\tilde{\phi}_{2}}$ below.
-$\begin{cases}\chi_{\phi_{1}}=\sum_{j,k}|j\rangle\langle k|\otimes\phi_{1}(|j\rangle\langle k|)/|A|,\\ \chi_{\phi_{2}}=\sum_{j,k}|j\rangle\langle k|\otimes\phi_{2}(|j\rangle\langle k|)/|A|,\\ \chi_{\tilde{\phi}_{1}}=\sum_{j,k}|j\rangle\langle k|\otimes\big{[}\phi_{1}(|j\rangle\langle k|)\oplus\bm{0}_{2}(|j\rangle\langle k|\big{]}/|A|,\\ \chi_{\tilde{\phi}_{2}}=\sum_{j,k}|j\rangle\langle k|\otimes\big{[}\bm{0}_{1}(|j\rangle\langle k|)\oplus\phi_{2}(|j\rangle\langle k|\big{]}/|A|.\end{cases}$
-Thus the following property holds.
+prove $\bm{(C_{4})}$.
+For any $\left\{\phi_{m}\right\}\subset\mathcal{C}_{AB}$ and any probability distribution $\left\{p_{m}\right\}$, we construct a auxiliary system $S$ with orthonormal basis $\left\{|m\rangle\right\}_{m}$ and define $\chi_{\phi}^{ABS}=\sum_{m}p_{m}\chi_{\phi_{m}}\otimes|m\rangle\langle m|$.
+Notably, we have:
+$\sum_{m}(I_{|A||B|}\otimes|0\rangle\langle m|)^{\dagger}(I_{|A||B|}\otimes|0\rangle\langle m|)=I_{|A||B||S|}$
+and $\Theta^{ABS}\in\mathcal{RSC}_{ABSA^{\prime}B^{\prime}S}$ with $\Theta^{ABS}(\cdot)=\sum_{m}(I_{|A||B|}\otimes|0\rangle\langle m|)\cdot(I_{|A||B|}\otimes|0\rangle\langle m|)^{\dagger}$. Then,
+$\Theta^{ABS}(\chi_{\phi}^{ABS})$
+$=$ $\sum_{m}(I_{|A||B|}\otimes|0\rangle\langle m|)\chi_{\phi}^{ABS}(I_{|A||B|}\otimes|0\rangle\langle m|)^{\dagger}$
+$=$ $\sum_{m}p_{m}\chi_{\phi_{m}}\otimes|0\rangle\langle 0|.$
+From $\bm{(C_{2})}$ and $\bm{(C_{5})}$, the following inequality holds:
+$\sum_{m}p_{m}C(\chi_{\phi_{m}})$
+$=$ $C(\chi_{\phi}^{ABS})$
+$\geqslant$ $C(\Theta^{ABS}(\chi_{\phi}^{ABS}))$
+$=$ $C(\sum_{m}p_{m}\chi_{\phi_{m}}),$
+that is, condition $\bm{(C_{4})}$ holds. ∎
+Therefore, from the above two theorems, we can give the Alternative framework for the imaginarity of quantum channels:
+Theorem 3.
+$\bm{(C_{3})}+\bm{(C_{4})}$ is equivalent to $\bm{(C_{5})}$ when condition $\bm{(C_{1})}$ and $\bm{(C_{2})}$ holds.
+When verifying whether a quantifier is an imaginarity measure, we can choose to verify either the conditions $\bm{(C_{1})}$$\bm{(C_{4})}$ or the conditions $\bm{(C_{1})}+\bm{(C_{2})}+\bm{(C_{5})}$.
 Property 4.
@@
 $C(p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}})=C(p_{1}\chi_{\tilde{\phi}_{1}}+p_{2}\chi_{\tilde{\phi}_{2}})=C(\chi_{\phi}).$ (3)
-The proof of Property 4 is obvious, here is a helpful example.
-Example 2.
-The following matrix transformations hold and they are invertible.
-$\begin{pmatrix}a_{11}&a_{12}&&&&\\ a_{21}&a_{22}&&&&\\ &&a_{33}&&&\\ &&&b_{11}&b_{12}&\\ &&&b_{21}&b_{22}&\\ &&&&&b_{33}\end{pmatrix}\xrightarrow{K_{1}(\cdot)K_{1}^{\dagger}}\begin{pmatrix}a_{11}&&a_{12}&&&\\ &a_{33}&&&&\\ a_{21}&&a_{22}&&&\\ &&&b_{22}&b_{21}&\\ &&&b_{12}&b_{11}&\\ &&&&&b_{33}\end{pmatrix}\xrightarrow{K_{2}(\cdot)K_{2}^{\dagger}}\begin{pmatrix}a_{11}&&a_{12}&&&\\ &b_{11}&&b_{12}&&\\ a_{21}&&a_{22}&&&\\ &b_{21}&&b_{22}&&\\ &&&&a_{33}&\\ &&&&&b_{33}\\ \end{pmatrix},$
-where
-$K_{1}=\begin{pmatrix}1&&&&&\\ &&1&&&\\ &1&&&&\\ &&&&1&\\ &&&1&&\\ &&&&&1\\ \end{pmatrix},K_{2}=\begin{pmatrix}1&&&&&\\ &&&&1&\\ &&1&&&\\ &&&1&&\\ &1&&&&\\ &&&&&1\\ \end{pmatrix}.$
-Due to $K^{\dagger}_{j}K_{j}=I$, then the invertible $\Theta_{j}\in\mathcal{RSC}$ with $\Theta_{j}(\cdot)=K_{j}\cdot K_{j}^{\dagger}$, $j=1,2$.
-Just like $\Theta_{1}$ and $\Theta_{2}$ in this example, we can construct multiple invertible channels such that
-$\Theta_{1}(\Theta_{2}(\cdots\Theta_{k}(p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}})\cdots))=p_{1}\chi_{\tilde{\phi}_{1}}+p_{2}\chi_{\tilde{\phi}_{2}}.$
-In this way, we can know that the equation (3) holds from the condition $\bm{(C_{2})}$.
+Proof.
+We employ the same idea as in Theorem 1, specifically using $\bm{(C_{2})}$ to establish equivalence.
+Without loss of generality, we assume $|A|=3$, $p_{1}\phi_{1}(|j\rangle\langle k|)\triangleq X_{jk}$ and $p_{2}\phi_{2}(|j\rangle\langle k|)\triangleq Y_{jk}$.
+Then,
+$p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}}=\begin{pmatrix}X_{11}&X_{12}&X_{13}&&&\\ X_{21}&X_{22}&X_{23}&&&\\ X_{31}&X_{32}&a_{33}&&&\\ &&&Y_{11}&Y_{12}&Y_{13}\\ &&&Y_{21}&Y_{22}&Y_{23}\\ &&&Y_{31}&Y_{32}&Y_{33}\end{pmatrix},$
+and
+$p_{1}\chi_{\tilde{\phi}_{1}}+p_{2}\chi_{\tilde{\phi}_{2}}=\begin{pmatrix}X_{11}&&X_{12}&&X_{13}&\\ &Y_{11}&&Y_{12}&&Y_{13}\\ X_{21}&&X_{22}&&X_{23}&\\ &Y_{21}&&Y_{22}&&Y_{23}\\ X_{31}&&X_{32}&&X_{33}&\\ &Y_{31}&&Y_{32}&&Y_{33}\end{pmatrix}.$
+In this way, we need to find some superchannels between $p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}}$ and $\chi_{\phi}$. We can set
+$K_{1}=\begin{pmatrix}I_{|B_{1}|}&&&&&\\ &&&I_{|B_{1}|}&&\\ &&&&I_{|B_{1}|}&\\ &I_{|B_{2}|}&&&&\\ &&I_{|B_{2}|}&&&\\ &&&&&I_{|B_{2}|}\end{pmatrix},$
+$K_{2}=\begin{pmatrix}I_{|B_{1}|}&&&&&\\ &I_{|B_{2}|}&&&&\\ &&I_{|B_{1}|}&&&\\ &&&I_{|B_{2}|}&&\\ &&&&I_{|B_{1}|}&\\ &&&&&I_{|B_{2}|}\end{pmatrix}.$
+It can be known that $K^{\dagger}_{j}K_{j}=I$, $j=1,2$, and the following transformations,
+$p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}}\xrightarrow{K_{1}(\cdot)K_{1}^{\dagger}}N\xrightarrow{K_{2}(\cdot)K_{2}^{\dagger}}p_{1}\chi_{\tilde{\phi}_{1}}+p_{2}\chi_{\tilde{\phi}_{2}},$
+where $N$ is
+$\begin{pmatrix}X_{11}&&&X_{12}&X_{13}&\\ &Y_{11}&Y_{12}&&&Y_{13}\\ &Y_{21}&Y_{22}&&&Y_{23}\\ X_{21}&&&X_{22}&X_{23}&\\ X_{31}&&&X_{32}&X_{33}&\\ &Y_{31}&Y_{32}&&&Y_{33}\end{pmatrix}.$
+We define $\Theta_{j}(\cdot)=K_{j}\cdot K_{j}^{\dagger}$, $j=1,2$. It is clear that $\Theta_{j}\in\mathcal{RSC}$. Then the inequality holds:
+$C(p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}})\geqslant C(p_{1}\chi_{\tilde{\phi}_{1}}+p_{2}\chi_{\tilde{\phi}_{2}}).$
+Since $K_{1}$ and $K_{2}$ are invertible, the inverse inequality also holds. Therefore, we have:
+$C(p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}})=C(p_{1}\chi_{\tilde{\phi}_{1}}+p_{2}\chi_{\tilde{\phi}_{2}}).$
+When the dimension $|A|$ takes other values, we can similarly construct many invertible real superchannels such that
+$\Theta_{1}(\cdots\Theta_{k}(p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}})\cdots)=p_{1}\chi_{\tilde{\phi}_{1}}+p_{2}\chi_{\tilde{\phi}_{2}}.$
+In this way, we can conclude the equation (3) holds. ∎
 4 Imaginarity measures of quantum channels
-The Choi theorem [27] tell us that the Choi matrix of the channel $\phi$ is positive semidefinite, so that $\chi_{\phi}\geqslant 0$, we can see that, for any quantum channel $\phi$ with $\phi(\cdot)=\sum_{n}K_{n}\cdot K_{n}^{\dagger}$, its Choi state $\chi_{\phi}=\sum_{j,k}|j\rangle\langle k|\otimes\phi(|j\rangle\langle k|)/|A|$, thus the trace:
-$\text{Tr}(\chi_{\phi})=\text{Tr}\Big{[}\sum_{j,k}|j\rangle\langle k|\otimes\phi(|j\rangle\langle k|)/|A|\Big{]}=\frac{1}{|A|}\sum_{j}\text{Tr}\Big{[}\phi(|j\rangle\langle j|)\Big{]}=\frac{1}{|A|}\sum_{j}1=1.$
-In this way, we can embed Choi states into quantum states:
-$\mathcal{C}_{AB}\hookrightarrow\mathcal{D}_{AB}.$
-It seems then that many imaginarity measures for quantum states can be equally applied to quantum channels.
-Thus in this section, we introduce three imaginarity measures of quantum channels, that is, measures via on the robustness, the trace norm and the sandwiched Rényi relative entropy, respectively.
-We can exactly verify these three measures in two ways.
-4.1 Measure via on the robustness
-The Measure via on the robustness is appeared in many resource theories [1, 18]. Similar to these, we define the imaginarity measure of quantum channels as follow.
+In this section, we introduce three imaginarity measures of quantum channels, that is, measures based on the robustness, the trace norm, and the sandwiched Rényi relative entropy, respectively.
+4.1 Measure based on the robustness
+The Measure based on the robustness has appeared in many resource theories [1, 18]. Similarly, we define the imaginarity measure of quantum channels as follows.
 $C_{r}(\phi)=\min_{\tilde{\phi}\in\mathcal{RC}_{AB}}\left\{s\geqslant 0:\frac{\chi_{\phi}+s\chi_{\tilde{\phi}}}{1+s}\triangleq\chi_{\psi},\psi\in\mathcal{RC}_{AB}\right\}$
@@
 Theorem 4.
-$C_{r}(\phi)$ is a proper imaginarity measure of quantum channels.
+$C_{r}(\phi)$ is an imaginarity measure of quantum channels.
 Before our proof, we need to introduce pseudomixture [27, 28, 29, 7] of $\chi_{\phi}$, that is,
@@
 Proof.
-We only verify $\bm{(C_{1})}$+$\bm{(C_{3})}$+$\bm{(C_{4})}$ since $\bm{(C_{2})}$ can be derived.
+We only verify $\bm{(C_{1})}$+$\bm{(C_{3})}$+$\bm{(C_{4})}$ since $\bm{(C_{2})}$ can be derived by $\bm{(C_{3})}$ and$\bm{(C_{4})}$.
 $\bm{(C_{1})}$:
-Obviously $C_{r}(\phi)\geqslant 0$. If $C_{r}(\phi)=0$, that means $\frac{\chi_{\phi}+0}{1+0}=\chi_{\psi}$, that is $\phi=\psi\in\mathcal{RC}_{AB}$. The converse is certainly true.
+It can be observed that $C_{r}(\phi)\geqslant 0$.
+When $C_{r}(\phi)=0$, this implies $\frac{\chi_{\phi}+0}{1+0}=\chi_{\psi}$, which means $\phi=\psi\in\mathcal{RC}_{AB}$.
+The converse is also clearly true.
 $\bm{(C_{3})}$:
-Let $\Theta\in\mathcal{RSC}_{ABA^{\prime}B^{\prime}}$ with $\Theta(\cdot)=\sum_{m}M_{m}\cdot M_{m}^{\dagger}$, for any quantum channel $\phi$ with its pseudomixture, under the requirement of $\bm{(C_{3})}$, $p_{m}=\text{Tr}(M_{m}\chi_{\phi}M_{m}^{\dagger})$ and $\chi_{\phi_{m}}=M_{m}\chi_{\phi}M_{m}^{\dagger}/p_{m}$.
-Then we obtain
-$M_{m}\chi_{\phi}M_{m}^{\dagger}=[1+C_{r}(\chi_{\phi})]M_{m}\chi_{\phi_{1}}M_{m}^{\dagger}-C_{r}(\chi_{\phi})M_{m}\chi_{\phi_{2}}M_{m}^{\dagger}.$
+Let $\Theta\in\mathcal{RSC}_{ABA^{\prime}B^{\prime}}$ with $\Theta(\cdot)=\sum_{m}M_{m}\cdot M_{m}^{\dagger}$.
+For any quantum channel $\phi$ with its pseudomixture, under the requirement of $\bm{(C_{3})}$, we have $p_{m}=\text{Tr}(M_{m}\chi_{\phi}M_{m}^{\dagger})$ and $\chi_{\phi_{m}}=M_{m}\chi_{\phi}M_{m}^{\dagger}/p_{m}$.
+Then, we obtain:
+$M_{m}\chi_{\phi}M_{m}^{\dagger}$
+$=$ $[1+C_{r}(\chi_{\phi})]M_{m}\chi_{\phi_{1}}M_{m}^{\dagger}-C_{r}(\chi_{\phi})M_{m}\chi_{\phi_{2}}M_{m}^{\dagger}.$
 Let
@@
 where $s_{m}=\frac{1}{p_{m}}C(\chi_{\phi})\text{Tr}(M_{m}\chi_{\phi_{1}}M_{m}^{\dagger})$.
-We get $\chi_{\phi_{m}}=(1+s_{m})\chi_{\phi_{1}}^{m}-s_{m}\chi_{\phi_{2}}^{m}$, that is, $\frac{\chi_{\phi_{m}}+s_{m}\chi_{\phi_{2}}^{m}}{1+s_{m}}=\chi_{\phi_{1}}^{m}$.
+We obtain $\chi_{\phi_{m}}=(1+s_{m})\chi_{\phi_{1}}^{m}-s_{m}\chi_{\phi_{2}}^{m}$.
+This means $\frac{\chi_{\phi_{m}}+s_{m}\chi_{\phi_{2}}^{m}}{1+s_{m}}=\chi_{\phi_{1}}^{m}$.
 Thus we get $C_{r}(\chi_{\phi_{m}})\leqslant s_{m}$ from definition, that is, $p_{m}C_{r}(\chi_{\phi_{m}})\leqslant C(\chi_{\phi})\text{Tr}(M_{m}\chi_{\phi_{1}}M_{m}^{\dagger})$.
-Therefore, we obtain
-$\sum_{m}p_{m}C_{r}(\chi_{\phi_{m}})$$\leqslant$ $\sum_{m}C(\chi_{\phi})\text{Tr}(M_{m}\chi_{\phi_{1}}M_{m}^{\dagger})$
+Therefore, we have:
+$\sum_{m}p_{m}C_{r}(\chi_{\phi_{m}})$
+$\leqslant$ $\sum_{m}C(\chi_{\phi})\text{Tr}(M_{m}\chi_{\phi_{1}}M_{m}^{\dagger})$
 $=$ $C(\chi_{\phi})\sum_{m}\text{Tr}(M_{m}^{\dagger}M_{m}\chi_{\phi_{1}})$
@@
 $\bm{(C_{4})}$:
-For simplicity, we only consider two quantum channels $\phi_{1}$ and $\phi_{2}$, and define $\phi=p\phi_{1}+(1-p)\phi_{2}$ with any $p\in(0,1)$, that is, $\chi_{\phi}=p\chi_{\phi_{1}}+(1-p)\chi_{\phi_{2}}$.
-The measurements $C_{r}$ for $\phi_{1}$ and $\phi_{2}$ are $C_{r}(\chi_{\phi_{1}})$ and $C_{r}(\chi_{\phi_{2}})$.
+For simplicity, we consider only two quantum channels $\phi_{1}$ and $\phi_{2}$, and define $\phi=p\phi_{1}+(1-p)\phi_{2}$ with any $p\in(0,1)$.
+This leads to $\chi_{\phi}=p\chi_{\phi_{1}}+(1-p)\chi_{\phi_{2}}$.
 Let $s_{p}=pC_{r}(\chi_{\phi_{1}})+(1-p)C_{r}(\chi_{\phi_{2}})$.
 Consider the pseudomixtures of $\chi_{\phi_{1}}$ and $\chi_{\phi_{2}}$:
-$\chi_{\phi_{1}}=[1+C_{r}(\chi_{\phi_{1}})]\chi_{\phi_{1,1}}-C_{r}(\chi_{\phi_{1}})\chi_{\phi_{1,2}}\mbox{ and }\chi_{\phi_{2}}=[1+C_{r}(\chi_{\phi_{2}})]\chi_{\phi_{2,1}}-C_{r}(\chi_{\phi_{2}})\chi_{\phi_{2,2}},$
+$\chi_{\phi_{1}}=[1+C_{r}(\chi_{\phi_{1}})]\chi_{\phi_{1,1}}-C_{r}(\chi_{\phi_{1}})\chi_{\phi_{1,2}},$
+$\chi_{\phi_{2}}=[1+C_{r}(\chi_{\phi_{2}})]\chi_{\phi_{2,1}}-C_{r}(\chi_{\phi_{2}})\chi_{\phi_{2,2}},$
 then let
 $\begin{cases}\chi_{\psi}=\frac{p[1+C_{r}(\chi_{\phi_{1}})]\chi_{\phi_{1,1}}+(1-p)[1+C_{r}(\chi_{\phi_{2}})]\chi_{\phi_{2,1}}}{1+s_{p}},\\ \chi_{\tilde{\phi}}=\frac{pC_{r}(\chi_{\phi_{1}})\chi_{\phi_{1,2}}+(1-p)C_{r}(\chi_{\phi_{2}})\chi_{\phi_{2,2}}}{s_{p}}.\end{cases}$
-We have $\chi_{\phi}=(1+s_{p})\chi_{\psi}-s_{p}\chi_{\tilde{\phi}}$, that is, $\chi_{\psi}=\frac{\chi_{\phi}+s_{p}\chi_{\tilde{\phi}}}{1+s_{p}}$, thus we get $s_{p}\geqslant C_{r}(\chi_{\phi})$.
+We have $\chi_{\phi}=(1+s_{p})\chi_{\psi}-s_{p}\chi_{\tilde{\phi}}$, that is, $\chi_{\psi}=\frac{\chi_{\phi}+s_{p}\chi_{\tilde{\phi}}}{1+s_{p}}$, thus we obtain $s_{p}\geqslant C_{r}(\chi_{\phi})$.
 Therefore, $pC_{r}(\chi_{\phi_{1}})+(1-p)C_{r}(\chi_{\phi_{2}})=s_{p}\geqslant C_{r}(\chi_{\phi})=C(p\chi_{\phi_{1}}+(1-p)\chi_{\phi_{2}})$. The condition $\bm{(C_{4})}$ holds. ∎
-4.2 Measure via on the trace norm
-To obtain the channel is real, we can consider the trace norm:
+4.2 Measure based on the trace norm
+The trace norm,
 $C_{t}(\phi)=||\chi_{\phi}-\chi_{\phi}^{*}||_{\text{Tr}}=\text{Tr}|\chi_{\phi}-\chi_{\phi}^{*}|.$
-where $\chi_{\phi}^{*}$ means the conjugate of $\chi_{\phi}$ and $|\chi_{\phi}-\chi_{\phi}^{*}|=\sqrt{(\chi_{\phi}-\chi_{\phi}^{*})^{\dagger}(\chi_{\phi}-\chi_{\phi}^{*})}$.
+can be used to construct a measure, where $\chi_{\phi}^{*}$ means the conjugate of $\chi_{\phi}$ and $|\chi_{\phi}-\chi_{\phi}^{*}|=\sqrt{(\chi_{\phi}-\chi_{\phi}^{*})^{\dagger}(\chi_{\phi}-\chi_{\phi}^{*})}$.
 Theorem 5.
@@
 The condition $\bm{(C_{1})}$ holds due to the definition of the trace norm.
-There is $||\chi_{\phi}-\chi_{\phi}^{*}||_{\text{Tr}}\geqslant||\Theta(\chi_{\phi})-\Theta(\chi_{\phi}^{*})||_{\text{Tr}}$ for any CPTP map [31], and it certainly holds for real superchannels.
-Since $\Theta$ is a real superchannel, we can know that $[\Theta(\chi_{\phi})]^{*}=\Theta(\chi_{\phi}^{*})$ due to
-$[\Theta(\chi_{\phi})]^{*}=\left[\sum_{m}M_{m}\chi_{\phi}M_{n}^{\dagger}\right]^{*}=\sum_{m}M_{m}\chi_{\phi}^{*}M_{n}^{\dagger}=\Theta(\chi_{\phi}^{*}).$
+Due to $||\chi_{\phi}-\chi_{\phi}^{*}||_{\text{Tr}}\geqslant||\Theta(\chi_{\phi})-\Theta(\chi_{\phi}^{*})||_{\text{Tr}}$ for any CPTP map [30], it certainly holds for real superchannels.
+Since $\Theta$ is a real superchannel, it can be seen that $[\Theta(\chi_{\phi})]^{*}=\Theta(\chi_{\phi}^{*})$ due to
+$[\Theta(\chi_{\phi})]^{*}$
+$=$ $\left[\sum_{m}M_{m}\chi_{\phi}M_{n}^{\dagger}\right]^{*}$
+$=$ $\sum_{m}M_{m}\chi_{\phi}^{*}M_{n}^{\dagger}$
+$=$ $\Theta(\chi_{\phi}^{*}).$
 Thus $\bm{(C_{2})}$ is satisfied.
 Under the requirement of $\bm{(C_{5})}$, from Property 4, we have:
-$C_{t}(\chi_{\phi})$ $=$ $C_{t}(p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}})=||(p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}})-(p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}})^{*}||_{\text{Tr}}$
-$=$ $||(p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}})-(p_{1}\chi_{\phi_{1}}^{*}\oplus p_{2}\chi_{\phi_{2}}^{*})||_{\text{Tr}}=||p_{1}(\chi_{\phi_{1}}-\chi_{\phi_{1}}^{*})\oplus p_{2}(\chi_{\phi_{2}}-\chi_{\phi_{2}}^{*})||_{\text{Tr}}$
-$=$ $p_{1}||(\chi_{\phi_{1}}-\chi_{\phi_{1}}^{*})||_{\text{Tr}}+p_{2}||(\chi_{\phi_{2}}-\chi_{\phi_{2}}^{*})||_{\text{Tr}}=p_{1}C_{t}(\chi_{\phi_{1}})+p_{2}C_{t}(\chi_{\phi_{2}}).$
+$C_{t}(\chi_{\phi})$
+$=$ $C_{t}(p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}})$
+$=$ $||(p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}})-(p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}})^{*}||_{\text{Tr}}$
+$=$ $||(p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}})-(p_{1}\chi_{\phi_{1}}^{*}\oplus p_{2}\chi_{\phi_{2}}^{*})||_{\text{Tr}}$
+$=$ $||p_{1}(\chi_{\phi_{1}}-\chi_{\phi_{1}}^{*})\oplus p_{2}(\chi_{\phi_{2}}-\chi_{\phi_{2}}^{*})||_{\text{Tr}}$
+$=$ $p_{1}||(\chi_{\phi_{1}}-\chi_{\phi_{1}}^{*})||_{\text{Tr}}+p_{2}||(\chi_{\phi_{2}}-\chi_{\phi_{2}}^{*})||_{\text{Tr}}$
+$=$ $p_{1}C_{t}(\chi_{\phi_{1}})+p_{2}C_{t}(\chi_{\phi_{2}}).$
 Therefore, $C_{r}(\phi)$ satisfies $\bm{(C_{1})}$+$\bm{(C_{2})}$+$\bm{(C_{5})}$. ∎
-4.3 Measure via on the sandwiched Rényi relative entropy
-The sandwiched Rényi relative entropy [32] defined for quantum states with the parameter $\alpha$ as follow:
+4.3 Measure based the sandwiched Rényi relative entropy
+The sandwiched Rényi relative entropy [31] is defined for quantum states with the parameter $\alpha$ as follows:
 $D_{\alpha}(\rho||\sigma)=\frac{1}{\alpha-1}\text{log }\text{Tr}(\sigma^{\frac{1-\alpha}{2\alpha}}\rho\sigma^{\frac{1-\alpha}{2\alpha}})^{\alpha},$
-where Now, we will introduce the imaginarity measure of quantum channels via on the sandwiched Rényi relative entropy, that is,
+Now, we will introduce the imaginarity measure of quantum channels based on the sandwiched Rényi relative entropy, defined as:
 $C_{\alpha}(\phi)=1-\text{Tr}\left[(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\chi_{\phi}(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\right]^{\alpha},$
-where $\frac{1}{2}\leqslant\alpha<1$. We have the following results.
+where $\frac{1}{2}\leqslant\alpha<1$. We have the following results:
 Theorem 6.
@@
 Proof.
-It is shown [32] that $D_{\alpha}(\rho||\sigma)\geqslant 0$ when $\frac{1}{2}\leqslant\alpha<1$ and equation holds if and only if $\rho=\sigma$, that is, $\text{Tr}(\sigma^{\frac{1-\alpha}{2\alpha}}\rho\sigma^{\frac{1-\alpha}{2\alpha}})^{\alpha}\leqslant 1$.
-Thus we consider Choi state as a quantum state and then we have, $C_{\alpha}(\phi)\geqslant 0$.
-The equation holds if and only if $\chi_{\phi}=\chi_{\phi}^{*}$.
-The condition $\bm{(C_{1})}$ is satisfied.
+It is shown [31] that $D_{\alpha}(\rho||\sigma)\geqslant 0$ when $\frac{1}{2}\leqslant\alpha<1$, and equation holds if and only if $\rho=\sigma$, which implies that $\text{Tr}(\sigma^{\frac{1-\alpha}{2\alpha}}\rho\sigma^{\frac{1-\alpha}{2\alpha}})^{\alpha}\leqslant 1$.
+Thus, considering Choi state as a quantum state, we find that $C_{\alpha}(\phi)\geqslant 0$.
+The equation holds if and only if $\chi_{\phi}=\chi_{\phi}^{*}$, satisfying the condition $\bm{(C_{1})}$.
 From [31], we have:
 $D_{\alpha}(\chi_{\phi}||\chi_{\phi}^{*})\geqslant D_{\alpha}\left[\Theta(\chi_{\phi})||\Theta(\chi_{\phi}^{*})\right]$
-$\iff$ log $\text{log }\text{Tr}\left[(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\chi_{\phi}(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\right]^{\alpha}\leqslant\text{log }\text{Tr}\left\{[\Theta(\chi_{\phi}^{*})]^{\frac{1-\alpha}{2\alpha}}\Theta(\chi_{\phi})[\Theta(\chi_{\phi}^{*})]^{\frac{1-\alpha}{2\alpha}}\right\}^{\alpha}$
-$\iff$ $\text{Tr}\left[(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\chi_{\phi}(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\right]^{\alpha}\leqslant\text{Tr}\left\{[\Theta(\chi_{\phi}^{*})]^{\frac{1-\alpha}{2\alpha}}\Theta(\chi_{\phi})[\Theta(\chi_{\phi}^{*})]^{\frac{1-\alpha}{2\alpha}}\right\}^{\alpha}$
-$\iff$ $\text{Tr}\left[(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\chi_{\phi}(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\right]^{\alpha}\leqslant\text{Tr}\left\{([\Theta(\chi_{\phi})]^{*})^{\frac{1-\alpha}{2\alpha}}\Theta(\chi_{\phi})([\Theta(\chi_{\phi})]^{*})^{\frac{1-\alpha}{2\alpha}}\right\}^{\alpha}$
+$\iff$ log $\text{log }\text{Tr}\left[(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\chi_{\phi}(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\right]^{\alpha}$
+$\leqslant\text{log }\text{Tr}\left\{[\Theta(\chi_{\phi}^{*})]^{\frac{1-\alpha}{2\alpha}}\Theta(\chi_{\phi})[\Theta(\chi_{\phi}^{*})]^{\frac{1-\alpha}{2\alpha}}\right\}^{\alpha}$
+$\iff$ $\text{Tr}\left[(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\chi_{\phi}(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\right]^{\alpha}$
+$\leqslant\text{Tr}\left\{[\Theta(\chi_{\phi}^{*})]^{\frac{1-\alpha}{2\alpha}}\Theta(\chi_{\phi})[\Theta(\chi_{\phi}^{*})]^{\frac{1-\alpha}{2\alpha}}\right\}^{\alpha}$
+$\iff$ $\text{Tr}\left[(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\chi_{\phi}(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\right]^{\alpha}$
+$\leqslant\text{Tr}\left\{([\Theta(\chi_{\phi})]^{*})^{\frac{1-\alpha}{2\alpha}}\Theta(\chi_{\phi})([\Theta(\chi_{\phi})]^{*})^{\frac{1-\alpha}{2\alpha}}\right\}^{\alpha}$
 $\iff$ $C_{\alpha}(\phi)\geqslant C_{\alpha}\left[\Theta(\phi)\right].$
@@
 Under the requirement of $\bm{(C_{5})}$, we have:
-$\text{Tr}\left[(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\chi_{\phi}(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\right]^{\alpha}$ $=$ $\text{Tr}\left\{\left[(p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}})^{*}\right]^{\frac{1-\alpha}{2\alpha}}(p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}})\left[(p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}})^{*}\right]^{\frac{1-\alpha}{2\alpha}}\right\}^{\alpha}$
+$\text{Tr}\left[(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\chi_{\phi}(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\right]^{\alpha}$
 $=$ $\text{Tr}\left\{\bigoplus_{j=1}^{2}p_{j}^{\frac{1-\alpha}{2\alpha}}p_{j}p_{j}^{\frac{1-\alpha}{2\alpha}}(\chi_{\phi_{j}}^{*})^{\frac{1-\alpha}{2\alpha}}\chi_{\phi_{j}}(\chi_{\phi_{j}}^{*})^{\frac{1-\alpha}{2\alpha}}\right\}^{\alpha}$
@@
 In this way, we obtain:
-$C_{\alpha}(\phi)=1-\text{Tr}\left[(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\chi_{\phi}(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\right]^{\alpha}=\sum_{j=1}^{2}p_{j}\left\{1-\text{Tr}\left[(\chi_{\phi_{j}}^{*})^{\frac{1-\alpha}{2\alpha}}\chi_{\phi_{j}}(\chi_{\phi_{j}}^{*})^{\frac{1-\alpha}{2\alpha}}\right]^{\alpha}\right\}=p_{1}C_{\alpha}(\phi_{1})+p_{2}C_{\alpha}(\phi_{2})$
+$C_{\alpha}(\phi)$
+$=$ $1-\text{Tr}\left[(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\chi_{\phi}(\chi_{\phi}^{*})^{\frac{1-\alpha}{2\alpha}}\right]^{\alpha}$
+$=$ $\sum_{j=1}^{2}p_{j}\left\{1-\text{Tr}\left[(\chi_{\phi_{j}}^{*})^{\frac{1-\alpha}{2\alpha}}\chi_{\phi_{j}}(\chi_{\phi_{j}}^{*})^{\frac{1-\alpha}{2\alpha}}\right]^{\alpha}\right\}$
+$=$ $p_{1}C_{\alpha}(\phi_{1})+p_{2}C_{\alpha}(\phi_{2})$
 The condition $\bm{(C_{5})}$ holds. ∎
 5 Conclusion
-We refine the framework of imaginarity of quantum channels which originally only has $\bm{(C_{1})}$ and $\bm{(C_{2})}$ for its measure.
-We add two conditions, strong monotonicity $\bm{(C_{3})}$ and convexity $\bm{(C_{4})}$.
+We introduce the framework of imaginarity of quantum channels, and introduce an alternative framework.
 Besides, we give some properties of imaginarity measure of quantum channels.
-Sometimes it is not easy to verify strong monotonicity and convexity, so we introduce an alternative framework, that is, $\bm{(C_{3})}+\bm{(C_{4})}$ is equivalent to $\bm{(C_{5})}$ when $\bm{(C_{1})}$ and $\bm{(C_{2})}$ hold.
-This is analogous to imaginarity or coherence of quantum states.
-Since Choi states can embed quantum states, we also think that imaginarity measures of quantum states can also be used for quantum channels and present three imaginarity measures for quantum channels, they are $C_{r}$, $C_{t}$ and $C_{\alpha}$ via on the robustness, the trace norm and the sandwiched Rényi relative entropy, respectively.
+We also give three imaginarity measures for quantum channels.
+They are $C_{r}$, $C_{t}$ and $C_{\alpha}$ based on the robustness, the trace norm and the sandwiched Rényi relative entropy, respectively.
 Acknowledgement